A Complete Bibliography of Publications in the IEEE symposia on Foundations of Computer Science

(FOCS) for 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 January 2018 Version 1.06

Title word cross-reference

#Knapsack [GKM^{+}11].

(1 + e) [GP13]. (1 + ϵ) [PW11a]. 1 [AALV11]. 1/2 [BFNS12]. 2 [BDSS11, CL12, DKN10, SCS11]. 3 [Cyg13, DKPS10, DMW13, GH11, GN13, Her11, KT12, SS10b]. 5 [BDD^{+}13]. _{#} [CLX10]. B [GS13a]. d [SS12b]. F [FLMS12, GS13a]. K [CV13b, COV13, ABS10b, BBMN11, Bel12, CV13a, CHK12, KT11, KK10, Ros10a]. M [DPV11]. m log n [KMP11]. O(c^{k}n) [BDD^{+}13]. O(log^{2} log n) [BG11]. O(log n)

[KSS13, BGS11]. O(log OPT ∗ OPT) [Rot13]. O(mn) [HKN13]. O(n^{2}) [PSSZ10]. P [CLP13b, LMS10]. q [KS12]. st [BSWN10].

-Approximate [GP13, PW11a, PW11b]. -Approximation [BDD^{+}13, BFNS12, GP10, KSS13]. -Centers [CHK12]. -Certificates [CLP13b]. -clique [Ros10a]. -Colorability [COV13]. -Colorable [DKPS10, KT12, KS12]. -cut [BSWN10]. -Deletion [FLMS12]. -Dimensional [Cyg13]. -Distinctness [Bel12]. -ellipsoid [DPV11]. -Factors [GS13a]. -Matching [GS13a]. -Means [ABS10b, KK10]. -Median [ABS10b]. -Node [CV13a]. -query [BDSS11].

-Regular [SS12b]. -SAT [GN13, Her11]. -Server [BBMN11, CV13b]. -way [KT11].

51st [IEE10]. 52nd [IEE11]. 53rd [IEE12]. 54th [IEE13].

Abstract [CLLM10]. AC0 [BIL12]. AC0-Circuits [BIL12]. Accelerated [LS13]. Access [SST11]. Active [BBBY12]. Adaptive [CLP10, SS13]. Adaptivity [IPW11]. Additive [BSLRZ12]. Advance [PSW13]. Adversarial [BGKS13, BK12]. Adversaries [GKM^{+}13]. Adversary [KMM11]. Affine [BSRZS12, HL13, Sha11]. Affine-Invariant [BSRZS12, HL13]. Against [KMM11, BK12, DHLAW10, GKM^{+}13, IV12]. Agnostically [Wim10]. Algebra [NN13]. Algebraic [FF11, FS13, GS13a]. Algorithm [AER13, BBMN11, Bel12, BDD^{+}13, BSW11, CDW13, CL12, FNS11, FW12, IPS13, JY11, iKRW11, KSS13, KK10, MS13, RT12, Vem10a, WZ13]. Algorithmic [CGS12, DR10b, Dwo11]. Algorithmics [Pru11]. Algorithms [AKO11, And10, ABS10a, Ban10, CLX10, CLLM10, CE11, CS13, CV13b, CCH^{+}12, DPV11, FLMS12, FR12, GS13a, GKMR11, GS12, HKK11, Le 12, LS13, Mad10, NN13, PS12, RW11, San10, SSY13]. All-or-Nothing [KK13]. All-Pairs [HKN13, PSSZ10]. Allocate [ABGG12]. Allocation [BFM13]. Almost [DKPS10, GT12a, HS12, KS12, Soh12]. Alphabets [Kop12]. Amortized [BG11, BR11]. Amplification [DLWZ11, Li12, Wee10]. Analysis [HR10, PS12]. Analyst [DNV12]. Analytic [Gop10]. Anarchy [LT10]. Annual [IEE10, IEE11, IEE12, IEE13]. Anticoncentrated [Kan12]. any [DPV11]. Applications [CGS12, HM10, JR11, Kan12, KL11, KLL^{+}12, Val12, Veg12, WN11]. Approach [AALV11, BSLRZ12, GSS11, Gop10, GLOV12, Vis12]. Approaching [KMP10]. Approximate [CFGS11, CLRS13, GP13, HKN13, LP10, PSW13, PW11a, PW11b, Vid13b, Vio11b]. Approximately [AFHH13]. Approximating [CV13a, DP10, GT12a, HPK13, LRV13, MS11a, MS11b, Rot13]. Approximation [AW13, AKO10, And10, AGS13, BG13, BDD^{+}13, BSW11, BFPS12, BFNS12, CLN13, CE11, CS13, CL12, Cyg13, FV13, FLMS12, FR12, GP10, GKMR11, GS11, JY11, KSS13, Mad10, SSY13]. Arbitrary [SS12a]. Area [AALV11]. Arithmetic [AIK11, GH11, GKKS13]. Art [Ano10c, Ano11d, Ano13d]. Arthur [HM10]. Aspects [HSS10]. Assembly [DPR^{+}10, DLP^{+}12]. Assumptions [CLP10]. Asymmetric [AKO10]. Asynchronously [ABGG12]. Attacks [DHLAW10]. Auction [LT10]. Auctions [Ala11, AFHH13, DV11, KSM10]. Augmented [CCKM10]. Author [Ano10a, Ano11a, Ano12a, Ano13a]. Available [Ano13g, Ano13h]. Average [FPRC13, KRT13]. Average-Case [KRT13]. Awards [Ano11b, Ano12b, Ano13b].

Back [Ano12c, Mad13]. Backdoors [GS13b]. Backward [WZ13]. Backyard [ANS10]. Bad [BKO11]. Balancing [SS12a]. Balls [CRSW11]. Bandits [BKS13, GKMR11]. Barrier [HKN13]. Based [Bel12, BDMI11, Mad10]. Baur [CGS12]. Bayes [LT10]. Bayesian [Ala11]. Beck [NNN12]. Becomes [CDW13]. Beliefs [CM11]. Berkeley [IEE13]. Better [AGS13, GMR^{+}12]. between [Mul12, WW10]. Beyond [BFPS12, AGM12, PR10]. Bias [HO11]. Bin [Rot13]. Bins [CRSW11, Rot13]. Bipartite [MSS13, RT12]. Bipartiteness [CMOS11]. Bit [CK11]. Black [BP12, DR10b, GLOV12, SS10b, Wee10]. Black-Box [DR10b, GLOV12, SS10b, Wee10]. Blackbox [Mul12]. Blindfolded [Vai11]. Bodlaender [EJT10]. Boosting [DRV10]. Bound [Alo10, ACNN11, BK10, BMMN11, BEO^{+}13, KL11, LP10, PR10, Rub13]. Bounded [Cyg13, DKN10, FR12, GS13b, JPY12, iKT11, KT11, KK13, SSY13]. Bounded-Round [JPY12]. Bounds [BIL12, BDSS11, CMT11, CK11, CS12, FV13, FPRC13, Her11, KLL^{+}12, KRT13, Lar12, PTW10, Pot10, RKS10, SS12a, SS10b]. Box [BP12, DR10b, GLOV12, SS10b, Wee10]. Branching [BRRY10, BV10, FS13]. Breaking [BEPS12, HKN13]. Broadcast [BFM13]. Brunswick [IEE12]. Bucket [BKKV10]. Budget [Sin10]. Buffer [AER13]. Bulk [GP10]. Buy [GP10]. Buy-at-Bulk [GP10]. Buyer [Ala11]. Buyers [Ala11]. Buys [Vio11b].

CA [IEE13]. California [IEE11]. Calls [HS12]. Candidate [GGH^{+}13]. Canonical [BCS^{+}13, Soh12]. Capacities [CHK12]. Capacity [GX13, KS11a]. Capture [CLX10]. Cascading [BEK^{+}11]. Case [ANS10, FPRC13, KRT13]. CC0 [LMS10]. Cell [Lar12]. Centers [CHK12]. Central [Mad13]. Certificates [CLP13b]. Chain [BK10]. Channels [GS10, KS11a]. Chaoses [Kan12]. Character [DLWZ11, KL11]. Chasing [COV13]. Chasm [GKKS13]. Choice [FF11]. Circuit [BSKK^{+}13, KLR12, Vio11a]. Circuit-SAT [BSKK^{+}13]. Circuits [AIK11, BIL12, CS12, GGH^{+}13, GH11, GKKS13, IPS13, SS10b]. Classes [iKR10]. Classical [Dru12]. clique [Ros10a, Ros10b]. Closed [GT12b, iKR10]. Clustering [ABS10b, CHLX13, GT12a, KK10]. Code [BGH^{+}12, DG13]. Codes [BSRZS12, BKS^{+}10, DGY10, GS10, GX13]. Coding [BKL11, BK12, CLL11, CPT13, GMS11, KL11]. Coin [BV10, HO11, JPY12, MPS10]. Collecting [KSS13]. Colorability [COV13]. Colorable [DKPS10, KT12, KS12]. Coloring [DG13, DMW13, KT12]. Column [BDMI11]. Column-Based [BDMI11]. Combinatorial [AALV11, Ala11, CVZ10, DV11, FW12, KT12, Lar11, LD11]. Combinatorics [BSLRZ12]. Commitment [CK11]. Commitments [GLOV12]. Committee [Ano10e, Ano10f, Ano11f, Ano11g, Ano12f, Ano12g, Ano13f, Ano13i]. Common [BP13]. Communication [BSLRZ12, BR11, BRWY13, GMS11, JPY12, KLL^{+}12, ST13]. commutative [FS13]. Commuting [AE11]. Competitive [AER13, BBMN11]. Complete [GN13]. Complexity [AALV11, AE11, AAGG11, AKO10, BSLRZ12, BSKK^{+}13, Ber12, BRWY13, CP11, GPS11, JPY12, KLL^{+}12, LMR^{+}11, LRV13, MPS10, Mul12, Ros10a, Ros10b, ST13, Vio10]. Composable [CLP10]. Compressibility [CS12]. Compression [Dru12]. Computation [CFGS11, Goy12, Wee10]. Computational [MPS10, Sly10, SS12b]. Computationally [GS10]. Compute [GR12]. Computer [IEE10, IEE11, IEE12, IEE13]. Computing [CDW12b, Mek12, Pru11, Vai11]. Concave [Veg12]. Concepts [Vem10b]. Concurrent [CLP13b]. Concurrently [Goy12]. Condition [RST^{+}11]. Conditions [AE11, iKRW11]. Configurations [SS10b]. Congestion [CFGS11, CL12, SCS11]. Conjecture [Sid13, TWXZ13, NNN12]. Connected [CV13a]. Connections [CLN13]. Connective [SSY13]. Connectivities [CLL11]. Connectivity [CNP^{+}11, FR12, Pot10]. Consequences [NNN12]. Constant [ANS10, BAETS10, BSKK^{+}13, BFM13, BSW11, BMMN11, CS12, CLP13b, HO11, Li13, SSY13, Vid13a]. Constant-Depth [CS12]. Constant-Round [CLP13b]. Constrained [DG13]. Constraint [CLRS13, FW12]. Construct [Zha12]. Constructing [GLOV12, HS12]. Constructions [GS10]. Constructive [Ban10, HSS10, LM12]. Contents [Ano10i, Ano11j, Ano12j, Ano13n]. Continual [BKKV10]. Continually [DLWW11]. Continuous [DHLAW10, FNS11, KS11a]. Continuous-Time [KS11a]. Contractions [CCH^{+}12]. Conversion [LMR^{+}11]. Convex [Vem10b]. Coordinate [LS13]. Copyright [Ano10b, Ano11c, Ano12d, Ano13c]. Correlated [GKMR11]. Correlation [BRS11]. Correlations [Val12]. Corrigendum [Vem10a]. Cost [AAZ10, BFM13, CCKM10, CV13a, RT12]. Costs [BGMS11]. Counterexample [NNN12]. Counting [GKM^{+}11, IW12, RST^{+}11, SS12b, Vio11b]. Coupled [BGKS13]. Coupled-Worlds [BGKS13]. Courcelle [EJT10]. Cover [Ano10c, Ano11d, Ano13d, Ano12c]. Coverings [DPV11]. Covers [KSM10]. Cryptographic [GM11, GLR10]. Cryptography [BKKV10, DHLAW10, GKM^{+}13]. CSP [CLX10]. CSPs [TZ12]. Cuckoo [ANS10]. Curvature [Sid13]. Cut [AGS13, CCH^{+}12, iKT11, KT11, Mad10, BSWN10]. Cut-Based [Mad10]. Cuts [KSM10]. Cutting [CVV12]. Cycles [BH13, CGS12, RW11].

D [AALV11, DMW13]. Data [BGKS13, HR10, JR11, LP10]. Decidable [CC10]. Deciding [DKT10]. Decision [BIL12, FKP11]. Decoding [BAETS10, Gop10]. Decompositions [And10]. Degree [DKN10, DG13, FR12, LMS10]. Degree- [DKN10]. Degree-Bounded [FR12]. Degrees [MSS13]. Delaunay [Rub13]. Delays [KS11a]. Deletion [FLMS12]. DeMorgan [KRT13]. Dense [CDK12]. Density [HPK12]. Dependent [CVZ10].

Depth [CS12, GH11, GKKS13, IPS13, SS10b, Vio11b]. Depth- [GH11, SS10b]. Derandomization [HKN13, Mul12]. Descent [LS13]. Design [AAZ10, BFM13, CDW12a, CDW13, CEGP10, CM11, DR10b, FR12, GKM^{+}13]. Designing [CCH^{+}12]. Designs [GK13]. Determinant [Bjö10]. Developments [Vai11]. Deviation [BIL12]. Devices [DLWW11]. Diagrams [CHLX13, HPK13]. Diameter [CGS12]. Dichotomy [CC10]. Differential [BBvdH12, BBDS12, DRV10, Dwo11, MMP^{+}10]. Differentially [BN10]. Digraphs [LNSWN12]. Dimension [ACNN11]. Dimensional [Cyg13, CDW12a]. Direct [BRWY13, Dru13, JPY12]. Directed [BH13, BKM^{+}11, CC10, CMPP13, Pot10]. Direction [IW12]. Dis [AAZ10]. Discrepancy [Ban10, Lar11, LM12]. Disjoint [And10, CL12, CMPP13, SCS11]. Disjointness [BP13, BEO^{+}13, ST13]. Disk [HLP13]. Dispersers [Sha11]. Dissection [AY10]. Distance [AKO10, GW12, HL13, PR10, PRT12]. Distinctness [BCM13, Bel12]. Distributed [BEPS12, FKP11]. Distribution [BS10]. Distributions [Vem10b, Vio10, Wim10]. Double [Tho13]. Down [HPK12]. Drawing [DMW13]. Driven [GKM^{+}13]. Dynamic [BGS11, GP13, HKN13, LP10]. Dynamics [Cha12].

Easy [Cha13]. Economics [AFHH13].

economies [AAZ10]. Edge

[And10, CL12, SCS11]. Edge-Disjoint [And10, CL12, SCS11]. Edges [LM12]. Edit [AKO10]. Efficient [BV11, BK12, CFGS11, DR10a, GMS11, GKL11, HM10, LS13, Wee10]. Efficiently [BWY12]. Eigenvalues [GS11]. Electrical [Mad13]. Element [BCM13]. ellipsoid [DPV11]. Embedding [Sid13]. Embeddings [NN13]. Encryption [BV11, GGH^{+}13, GH11, Vai11]. Entangled [IV12, Vid13b]. Entropy [Li13, Sha11]. Enumerative [DPV11]. Environmentally [CLP13a]. eps [PW11b]. Epsilon [Alo10]. Epsilon-Nets [Alo10]. Equal [ST13]. Equals [BR11]. Equilibria [CFGS11, Veg12]. Equilibrium [GPS11]. Equivalence [Mul12]. Equivalences [RW11, WW10]. Errors [KLR12]. Estimating [HL13, SS10a]. Estimation [HPK12]. Estimators [VV11]. Euclidean [BG13]. Euler [CS13]. Evaluating [Lar12]. Evaluation [CRSW11]. Everywhere [CDK12]. Everywhere-Sparse [CDK12]. Evolution [Kan11a]. Exchange [CVZ10]. Exclusion [BG11]. Exists [ST13]. Exists-Equal [ST13]. Expander [CLL11]. Expanders [Tho13]. Expanding [Ala11]. Expansion [BFK^{+}11, GT12a, LRV13, PTW10]. Expected [LD11]. Explicit [GMS11, GS10, GK13]. Exploiting [BGKS13]. Exponential [CNP^{+}11, HK12, BK10]. Exponentially [ES11]. Extendability [MM10]. Extension [KL11, MM10]. Extensions [HLP13]. Extractors [DLWZ11, Li12, Li13, Vio11a]. Extreme [CD11]. Extreme-Value [CD11]. Eyes [GT12b].

Factor [BSW11]. Factors [GS13a]. Failures [BEK^{+}11]. Fair [BFM13]. Families [BS10, CRSW11, MSS13]. Fast [GW12, Mad10, Tho13, WY10]. Faster [BCS^{+}13, CRSW11, GS12, Her11, Le 12, LS13, NN13]. Fault [DPR^{+}10]. Fault-Tolerance [DPR^{+}10]. Feasible [Sin10]. Fields [BDSS11, HSS11, LMS10]. Fighting [San10]. Finding [Ber12, DKPS10, KS12, Val12]. Finite [BDSS11, LMS10]. First [DKT10]. First-Order [DKT10]. Fixed [CMPP13, iKT11, KT11]. Fixed-Parameter [CMPP13, iKT11, KT11]. Flipping [HO11, MPS10]. Flow [BSW11, BKM^{+}11, KK13]. Flows [KSM10, LNSWN12, Mad13, She13, Veg12]. Fools [DKN10]. Forest [HLP13]. Foreword [Ano10d, Ano11e, Ano12e, Ano13e]. Forming [BKO11]. Forms [BCS^{+}13]. Formula [KRT13, San10]. Formulas [BPW11, GKL11, KLR12]. Forward [WZ13]. Forward-Backward [WZ13]. Foundations [IEE10, IEE11, IEE12, IEE13]. Fourier [Gop10, LMS10, TWXZ13]. Fourier-Analytic [Gop10]. FPT [CCH^{+}12, FLMS12]. FPTAS [GKM^{+}11]. Fractionality [KK13]. Framework [BGS10, HS13]. Free [CV13a, WN11]. Frequency [BCM13]. Friendly [CLP13a]. Frugal [CEGP10, KSM10]. Fully [BGS11, BV11, GH11, GP13, Vai11]. Functional [GGH^{+}13]. Functions [BWY12, COPV13, DKN10, FV13, HO11, JR11, Kan11b, Kan12, Zha12]. Fuzzy [DPR^{+}10].

Gallai [SS10b]. Game [KS11b]. Games [ABS10a, CFGS11, GH13, HM10, KMM11, Vid13b]. Gap [GX13]. Garble [AIK11]. Gaussian [Kan12, Mek12, Vem10b]. Gaussians [Kan11b, MV10]. General [Her11, Kop12]. Generalized [HPK13, HKK11, LT10, Veg12]. Generator [HS12]. Generators [BRRY10, GMR^{+}12, LMS10]. Genus [CS13]. Geometric [Mul12, RKS10]. Geometry [BP10, IKM10]. Gibbard [IKM10]. Global [BRS11]. Going [AGM12]. Goldreich [TW11]. Graph [BFK^{+}11, Bel12, CC10, CLL11, DMW13, GT12a, GS11, iKRW11]. Graphic [MS11a, MS11b]. Graphs [BCS^{+}13, BSWN10, BKM^{+}11, BDR^{+}12, CV13a, CLL11, CMOS11, DFP10, DKPS10, DKT10, KT12, KK13, KS12, Mad10, MSS13, PRT12, RT12, Ros10a, Ros10b, SCS11, SSY13, SS12b, WN11]. Greedy [FNS11, PS12]. Green [Pru11]. Grid [RST^{+}11]. Grothendieck [BMMN11]. Group [CDW12b, Lar11].

Halfspaces [Vem10a]. Hamiltonian [BH13]. Hamiltonians [AE11]. Hamiltonicity [Bjö10]. Hard [CHK12, Vid13b]. Hardness [CLP10, DKPS10, DG13, KS12, SS12b]. Hardnesses [CLN13]. Hash [CRSW11]. Hashing [ANS10]. Hierarchies [BRS11, BFPS12]. Hierarchy [GS11, GS12]. High [PSSZ10, Tho13]. Higher [GS11, Lar12]. Hold [Her11]. Hole [BKKV10, HPK12]. Holographic [CLX10]. Homomorphic [BV11, GH11, Vai11]. Homomorphisms [CC10]. Homotopy [GPS11, LR12]. Howson [GPS11]. Hyatt [IEE12]. Hypergraph [DG13].

Identification [WY12]. Identity [FS13, Mul12, SS10b]. IEEE [IEE10, IEE11, IEE12, IEE13]. II [AMS12]. iii [Ano10k, Ano11l, Ano13p]. Images [TR10]. Imperfect [RT12]. Implies [HO11]. Impossibility [BP12, BN10, FF11]. Improved [AER13, Cyg13, GW12, KRT13, RST^{+}11, RKS10, San10, SS10b]. Incentive [GKM^{+}13]. Incentive-Driven [GKM^{+}13]. Incentives [CDW13]. Increasing [SS10a]. Independence [CV13a, DKN10, Tho13]. Independence-Free [CV13a]. Independent [AW13, CLN13, DDO^{+}13, DKPS10, KS12, Li13, RST^{+}11]. Index [Ano13a, CCKM10, Ano10a, Ano11a, Ano12a]. Indexing [Kop12]. Indistinguishability [GGH^{+}13]. Induced [CLN13, CHLX13]. Infinity [PRT12]. Influence [Cha12]. Information [Ano11h, Ano13j, BP13, BR11, CCKM10, KLL^{+}12, Ano12h]. Insecurity [LW10]. Instance [Dru12]. Integer [DDO^{+}13, GS11]. Interactive [BK12, CS12, CPT13, GMS11, IV12]. Interlacing [MSS13]. Intersection [Vem10a]. Intrinsically [DLP^{+}12]. Invariant [BSRZS12, BGS10, HL13, Wim10]. Inverse [Vid13a]. Irrelevant [KW12]. Isomorphism [BWY12]. Isomorphism-Testable [BWY12]. Iterative [FR12, LMP13]. Itself [BBDS12].

Jersey [IEE12]. Johnson [BBDS12]. Join [BDR^{+}12]. Juntas [FV13, Val12].

k-clique [Ros10b]. K-Vertex-Disjoint [CMPP13]. k-way [iKT11]. Kernelization [FLMS12, KW12]. Key [BKKV10]. Kinetic [Rub13]. Klee [Cha13]. Knapsacks [BKS13, GKMR11]. Knowing [PSW13]. Knowledge [CLP13b, CPT13]. Knowledge-Preserving [CPT13]. Krivine [BMMN11].

L1 [ACNN11]. Language [CCKM10]. Large [BIL12, CLRS13]. Lasserre [GS11]. Lattice [DPV11]. Law [AALV11]. Layered [DMW13]. Layout [CKR11]. Layouts [DMW13]. LCCs [BDSS11]. Leakage [BKKV10, GR12, LW10]. Leaky [DLWW11]. Learnability [MV10]. Learning [AGM12, BS10, Bel12, CHW10, DDO^{+}13, Val12, Vem10a, Vem10b, Wim10]. Learning-Graph-Based [Bel12]. Least [BEK^{+}11]. Lemke [GPS11]. Lemma [HSS10, Mul12]. Levin [TW11]. Lexicographic [BKL11]. Lie [CDW12b]. Light [ES11]. Lighter [ES11]. Limitations [DV11]. Limits [BFPS12, Dru12, MMP^{+}10]. Lindenstrauss [BBDS12]. Line [BK10, Kop12]. Linear [AY10, ACNN11, BG13, BSRZS12, BBvdH12, BGS10, BSWN10, BKM^{+}11, BFPS12, BFNS12, DP10, HS12, KMP10, KMP11, LS13, NN13, She13, TZ12, VV11, Alo10, BKL11, GH13]. Linear-Invariant [BGS10]. Links [Ano13k, Ano13l]. Lipschitz [JR11, MM10]. List [BAETS10, Kop12]. LMP [KSS13].

Load [SS12a]. Local [AE11, BAETS10, Cyg13, DJW13, FKP11, GT12a, GS12, HSS10]. Locality [BEPS12]. Locally [BSRZS12]. Log [TWXZ13, Vid13a]. Log-Rank [TWXZ13]. Logspace [EJT10]. Long [BGH^{+}12, DG13]. Longest [SS10a]. Lossy [MS13]. Lovász [HSS10]. Low [DG13, LMS10]. Low-Degree [DG13, LMS10]. Lower [Alo10, ACNN11, BIL12, BDSS11, CS12, FPRC13, KLL^{+}12, KRT13, Lar12, LP10, PTW10]. LP [CLRS13, CHK12]. LTCs [Vid13a]. LWE [BV11].

Machine [CHW10]. Made [Cha13]. Making [BGH^{+}12]. malleability [Wee10]. malleable [DLWZ11, GLOV12, Li12]. Management [AER13]. Manipulation [IKM10]. Many [Ala11]. Market [Veg12]. Markov [CKR11]. martingale [GKMR11]. Matchgates [CLX10]. Matching [BGS11, CLN13, Cyg13, DP10, DGY10, GS13a, GT12b, MP12, PS12]. Matchings [CVV12, CGS12, GP13, IW12, Mad13, MS11a, MS11b, RT12]. Matrix [BDMI11, Le 12, WY10, WW10]. Matroid [FW12]. Max [LNSWN12, BFK^{+}11]. Maximal [BGS11]. Maximization [BGMS11, BFNS12, CDW12a, FNS11, FW12]. Maximizing [LD11]. Maximum [AW13, BKM^{+}11, DP10, PS12, SCS11, She13]. Means [ABS10b, KK10]. Measure [Cha13]. Mechanism [CDW12a, CDW13, CEGP10, CM11, DR10b, HR10, HK12]. Mechanisms [Ala11, BN10, DV11, Sin10]. Median [ABS10b]. Medium [SST11]. Membership [LP10]. Memory [BKKV10, DHLAW10]. Memoryless [CV13b]. Merlin [HM10]. Message [BEO^{+}13]. Message-Passing [BEO^{+}13]. Method [CVV12, GPS11]. Methods [LS13]. Metric [MM10, PTW10]. Milder [GMR^{+}12]. Min [BFK^{+}11, BSWN10, Li13, RT12]. Min-Cost [RT12]. Min-Entropy [Li13]. Min-max [BFK^{+}11]. Minimax [DJW13]. Minimization [Ban10, HPK13, LM12]. Minimum [AAZ10, CV13a, iKT11, KT11, RW11]. Minimum-Cost [AAZ10, CV13a]. Minor [iKR10, iKRW11, WN11]. Minor-Closed [iKR10]. Minor-Free [WN11]. Mixing [CMT11, RST^{+}11, SSY13]. Mixtures [MV10]. Model [BEO^{+}13, CLP10, DLP^{+}12, Goy12, Lar11]. Models [AGM12, KMM11, SS12b]. Moments [BCM13]. Monotone [FPRC13, Pot10, Ros10a, Ros10b]. Moser [HS13]. Motions [Rub13]. Muller [BKS^{+}10, Gop10]. Multi [CDW12a, IV12]. Multi-dimensional [CDW12a]. Multi-prover [IV12]. Multicommodity [KK13]. Multidimensional [CD11]. Multilinear [GKL11]. Multiple [BKM^{+}11]. Multiple-Sink [BKM^{+}11]. Multiple-Source [BKM^{+}11]. Multiplication [BBvdH12, Le 12, WY10]. Multiplicative [HR10]. Multiplicities [CDW12b]. Multivariate [HSS11]. Multiway [CE11]. Mutual [BG11].

Narrow [Ber12]. Nash [CFGS11, LT10]. Navigating [Mad13]. Near [ACNN11, BSWN10, BKM^{+}11, BDMI11, DP10, PTW10, Rub13]. Near-Linear [BSWN10, BKM^{+}11, DP10]. Nearly [HK12, KMP11, She13]. Nearly- [KMP11]. negative [CC10]. Neighbor [PTW10]. Nested [AY10]. Nets [Alo10]. Network [AAZ10, BFM13, BKL11, CLL11, FR12, SS12a]. Networks [BEK^{+}11, FPRC13, Pot10, SS13]. Nevada [IEE10]. NEXP [IV12]. Node [CV13a, FR12, HLP13, KSS13, NPS11]. Node-Connectivity [FR12]. Node-Weighted [HLP13, NPS11]. Noether [Mul12]. Noise [BK12]. Non [Alo10, BP12, BKL11, CC10, CHK12, DLWZ11, FS13, GH13, GLOV12, GKMR11, Li12, Sid13, Wee10]. Non-Black-Box [BP12]. Non-commutative [FS13]. Non-linear [Alo10, BKL11, GH13]. Non-malleability [Wee10]. Non-malleable [DLWZ11, GLOV12, Li12]. Non-martingale [GKMR11]. Non-negative [CC10]. Non-positive [Sid13]. Non-uniform [CHK12]. Nondeterministic [Dru13]. Nonrepetitive [DMW13]. Norm [DPV11, KK10, TWXZ13]. Normalization [Mul12]. Nothing [KK13]. notice [Ano12d]. NP [Vid13b]. NP-Hard [Vid13b]. Number [BAETS10, DFP10, HS12, Li13]. Numerical [NN13].

O [GP10]. Obfuscation [BP12, GGH^{+}13]. Objectives [GS11]. Oblivious [FS13]. October [IEE10, IEE11, IEE12, IEE13]. On-Line [BK10, Kop12]. Once [BPW11, FS13]. One [COPV13, GP10, HO11, KS11b]. One-Way [COPV13, HO11]. Ones [ES11]. Online [HLP13, MP12, NPS11]. Operations [ANS10]. Operators [BBvdH12, MM10]. Opinion [BKO11]. Optimal [AFHH13, BKS^{+}10, BDMI11, BN10, CD11, CDW12a, CK11, FV13, FLMS12, GT12a, GS10, HSS11, HK12, Sid10, Soh12, BBvdH12]. Optimality [KMP10]. Optimization [CHW10, LD11]. Oracle [BSWN10]. Oracles [GH13, GW12, PR10, PRT12]. Order [AE11, DKT10]. Ordered [Kop12]. Organizing [Ano10e, Ano11f, Ano12f, Ano13f]. OSNAP [NN13]. Our [GT12b]. Overcoming [BKKV10]. Own [BKO11].

Packing [Rot13]. Page [Ano10b, Ano10h, Ano10j, Ano10k, Ano11c, Ano11k, Ano11l, Ano13c, Ano13o, Ano13p, Ano12k, Ano12l]. Paintings [HLP13]. Pairs [HKN13, PSSZ10]. Palm [IEE11]. Parallel [CP11, JY11, LW10]. Parameter [CMPP13, iKT11, KT11]. Parameterized [CNP^{+}11]. Parities [Val12]. Parity [BH13, iKRW11]. Partial [HS13, WY12]. Partially [BWY12]. Partition [CE11]. Partitioning [BFK^{+}11, BK10, GS11]. Partitions [BDR^{+}12]. Party [JPY12, MMP^{+}10]. Passing [BEO^{+}13]. Path [Mad13, WW10]. Paths [And10, BSW11, CL12, CMPP13, GS13a, GW12, HKN13, PSSZ10, SCS11, WY10, WZ13]. Pathwidth [Cyg13]. PCA [Vem10b]. PCPs [BSKK^{+}13, DG13]. Column [DR10a]. Perebor [San10]. Perfect [CVV12, IW12]. Permanent [Vis12]. Permutation [Wim10]. Permutations [NNN12, RKS10]. Plain [CLP10, Goy12]. Planar [Alo10, BSWN10, BKM^{+}11, CMPP13, CMOS11, DFP10, FLMS12, KK13, LNSWN12, SCS11, Sid13, CLX10]. Planarization [Sid10]. Plane [CVV12]. Play [KMM11]. Player [Vid13b]. Playing [GH13]. Polar [GX13]. Polarization [GX13]. Poly [Vid13a]. Poly-Log [Vid13a]. Polylogarithmic [AKO10, BBMN11, CL12, Li13, SS10a]. Polylogarithmic-Competitive [BBMN11]. Polynomial [BS10, CVV12, GX13, Kan11b, Kan12, MV10, MS13, Mul12, Sha11]. Polynomials [HSS11, Lar12, LMS10]. Poorly [Kan12]. Population [MS13, WY12]. Positive [Goy12, JY11, Sid13]. Power [BKL11, DNV12, IPW11, TZ12, VV11]. PPSZ [Her11]. Precisely [CLX10]. Precision [AKO11]. Predecessor [Kop12]. Preprocessing [BSWN10]. Presence [GR12]. Preserves [BBDS12]. Preserving [CPT13, HR10]. PRG [Kan11b]. Price [BFM13, LT10]. Pricing [CD11, CLN13]. Prime [HSS11]. Privacy [BGKS13, BBDS12, DLWZ11, DJW13, DRV10, Dwo11, DNV12, HR10, JR11, Li12, MMP^{+}10]. Privacy-Preserving [HR10]. Private [BN10, HK12]. Prize [KSS13]. Probability [Dru13, PSSZ10]. Probe [Lar12]. Problem [And10, BBMN11, BV10, Cha13, CV13b, CMPP13, GSS11, HKK11, KK13, PS12, Vis12]. Problems [ABS10a, CS13, CCH^{+}12, CNP^{+}11, GKM^{+}11, LD11, Mad10, NPS11, ST13, WW10]. Proceedings [Ano13g, Ano13h, IEE10, IEE11, IEE12, IEE13]. Processes [Mek12]. Product [Dru13, HM10, JPY12]. Production [BGMS11]. Products [BKL11, BRWY13]. Profile [GT12a]. Profit [BGMS11]. Program [Ano10f, Ano11g, Ano12g, Ano13i]. Programming [BRS11, GS11, JY11, TZ12]. Programs [BFPS12, BRRY10, BV10, FS13]. Promise [Dwo11]. Proof [FF11, IKM10, IV12]. Proofs [Ber12]. Properties [BGS10, CVZ10, DKT10, HL13, TR10]. Property [BBBY12, Soh12]. Protocol [GKM^{+}13]. Protocols [CLP13a, GM11, GLR10, KLL^{+}12]. Prover [KS11b, IV12]. Provers [IV12]. Proximity [HPK12, HPK13]. PSD [GS11]. Pseudorandom [BRRY10, GMR^{+}12, HS12, LMS10]. Pseudorandomness [BPW11, BV10, IMZ12]. PTAS [ABS10b, Mek12]. Public [BKKV10, JPY12]. Public-Coin [JPY12]. Public-Key [BKKV10]. Publisher [Ano12h, Ano11h, Ano13j]. Pure [CFGS11, LT10].

QBF [San10]. QMA1 [GN13]. QMA1-Complete [GN13]. Quadratic [GS11, Rub13, TW11]. Quantitative [IKM10]. Quantum [AALV11, Bel12, CK11, Dru12, GN13, HM10, LMR^{+}11, Zha12]. Quasi [BBvdH12]. Quasi-optimal [BBvdH12]. Quasipolynomial [FS13]. Quasipolynomial-Time [FS13]. Queries [BAETS10, Kop12]. Query [AKO10, BSKK^{+}13, LMR^{+}11, BDSS11]. Queue [DFP10, DMW13]. Queues [SST11]. Quick [Ano13k, Ano13l].

Rabbit [HPK12]. Raecke [And10]. Ramanujan [MSS13]. Random [CMT11, CMOS11, DDO^{+}13, GKL11, Ros10a, Ros10b, Vem10a, Zha12, KMM11]. Randomized [CVZ10, CV13b, CCH^{+}12, DR10b, DV11, GSS11, PS12, SS12a]. Randomness [CP11, Vio11b]. Range [AMS12, Lar11]. Rank [BSLRZ12, SS10b, TWXZ13]. Rarity [Lev12]. Rate [BSKK^{+}13, GS10, Vid13a]. Rates [DJW13]. Rational [GKM^{+}13]. Rationality [GLR10]. Read [BPW11, FS13]. Read-Once [BPW11, FS13]. Recognition [CCKM10]. Recombination [Kan11a]. Reconstruction [BDMI11, GKL11, JR11]. Recovery [IPW11, MS13, PW11a, PW11b, WY12]. Rectangles [AW13]. Rectangular [Le 12]. Reducing [CDW12a]. Reduction [ACNN11]. Reductions [Dru13, DR10b]. Reed [BKS^{+}10, Gop10]. Regency [IEE12]. Regular [BCS^{+}13, BRRY10, SS12b]. Related [ABS10a, CS13, GKM^{+}11, NPS11]. Relating [BSLRZ12]. Relaxations [CLRS13]. Renaming [AAGG11]. Reordering [AER13]. Repetition [CP11, LW10]. Replacement [GW12, WY10]. Representation [ANS10]. Representations [CDW12b]. Representative [KW12]. Requires [CLRS13, HS12]. Resampling [HS13]. Resettability [COPV13]. Resilience [LW10]. Resilient [BKKV10, KLR12]. Restrictions [GMR^{+}12]. Results [Goy12]. Revenue [CDW12a]. Reviewers [Ano10g, Ano11i, Ano12i, Ano13m]. Rewards [MP12]. Robust [FF11, HPK12]. Roster [Ano10h]. Round [CLP13b, JPY12, KS11b, Wee10].

Round-Efficient [Wee10]. Rounding [BRS11, CLLM10, CVZ10, CHK12, FR12, GSS11, GS12]. Row [DR10a, LMP13]. Row/Column [DR10a].

Salesman [GSS11, Vis12]. Sampling [AKO11, BIL12, CMT11, DR10a, LMP13, RST^{+}11, Vem10a]. SAT [BSKK^{+}13, Dru13, GS13b, GN13, Her11]. Satisfaction [CLRS13]. Satisfiability [IPS13, San10, Soh12]. Satterthwaite [IKM10]. Scale [AAZ10]. Scaling [RT12]. Scheduling [BP10]. Scheme [BG13]. Schemes [AW13, GS11]. Science [IEE10, IEE11, IEE12, IEE13]. SDD [KMP10, KMP11]. SDP [GS12]. Search [Cyg13, HPK12, HPK13, PTW10]. Searching [AMS12, Lar11]. Second [LT10]. Secrets [DLWW11]. Secure [Goy12, Wee10]. Security [CLP10]. Seeding [SS13]. Selection [DR10a, GPS11]. Self [DPR^{+}10]. Self-Assembly [DPR^{+}10]. Semi [KMM11]. Semi-random [KMM11]. Semialgebraic [AMS12]. Semidefinite [BRS11, JY11]. Semimeasures [Lev12]. Sensitivity [GW12]. Separator [iKR10, WN11]. Separators [DMW13]. Sequence [SS10a]. Sequential [GLR10]. Server [BBMN11, CV13b]. Set [AW13, BFK^{+}11, BEO^{+}13, CLN13, CM11, PSW13, ST13]. Set-Theoretic [CM11]. Sets [AMS12, DKPS10, KS12, KW12, RST^{+}11]. Settling [MV10]. Shallow [ES11, WN11]. Shallow-Light [ES11]. Sharp [CMT11]. Short [KLR12]. Short-Circuit [KLR12]. Shorter [BGH^{+}12]. Shortest [CGS12, GS13a, HKN13, PSSZ10, WZ13]. Shrinkage [IMZ12]. Simple [AFHH13, GS10, Tho13]. Simpler [Her11]. Simulation [BP12]. Simultaneous [COPV13, GP10]. Single [Ala11, CNP^{+}11, GP10, GW12, LNSWN12, WZ13]. Single-Sink [GP10]. Single-Source [GW12, WZ13]. Sink [BKM^{+}11, GP10]. Sinks [LNSWN12]. Size [iKT11, KT11, KRT13, PSW13]. Sliding [BCM13]. Small [BFK^{+}11, HSS11, Kan11b]. Smaller [BMMN11, CRSW11]. Social [FF11, HK12, SS13]. Solutions [GPS11]. Solver [KMP11]. Solvers [Dru13, GS12]. Solving [AY10, CNP^{+}11, KMP10, LS13]. Some [NNN12]. Sorting [HKK11]. Sound [IV12]. Soundness [KS11b, Vid13a]. Source [BKM^{+}11, GW12, LNSWN12, Li12, WZ13]. Sources [Li13, Sha11, Vio11a]. Space [HPK12]. Spanners [CDK12]. Spanning [ES11]. Sparse [BSRZS12, CDK12, DKT10, IPS13, IPW11, PRT12, PW11a, PW11b, ST13, TR10]. Sparser [NN13]. Sparsest [AGS13]. Sparsifiers [CLLM10, MM10]. Sparsity [TWXZ13]. Spatial [SSY13]. Spectral [CEGP10, KK10, TWXZ13]. Spectrum [LMS10]. Speed [GX13, Rub13]. Spin [SS12b]. Split [BDR^{+}12]. Sponsors [Ano13f]. Springs [IEE11]. Squashing [GH11]. Stability [ABS10b, BFM13]. Standard [BV11, CLP10]. State [DNV12, LMR^{+}11]. Stateless [GM11]. States [AALV11, HM10]. Statistical [BGKS13, DJW13]. Steiner [BDR^{+}12, ES11, HLP13, KSS13, NPS11]. Stochastic [LD11, MP12, Sid10]. Storing [DLWW11]. Strassen [CGS12]. Streaming [AKO11, CCKM10]. Strictly [BMMN11]. Strong [BDR^{+}12, DPR^{+}10, GS13b, KS11b, Vid13a]. Strongly [BCS^{+}13]. Structure [Kan12]. Structures [CVZ10, LP10]. Study [Kan12, KMM11]. Sub [BK10, Sha11]. Sub-exponential [BK10]. Sub-polynomial [Sha11]. Subcubic [WW10]. Subexponential [ABS10a, CLN13]. Subgraphs [CV13a, CDK12]. Subject [FW12]. Sublinear [BSKK^{+}13, CHW10, HPK12]. Submodular [BFNS12, CE11, FNS11, FV13, FW12]. Subquadratic [Val12]. Subset [DR10a]. Subsets [Kop12]. Subspace [GK13, NN13]. Success [Dru13]. Succinct [ANS10]. Such [AE11]. Suffices [GP10]. Sums [Bjö10, DDO^{+}13, DLWZ11, KL11]. Supremum [Mek12]. Surfaces [CMT11, LR12]. Susceptible [BEK^{+}11]. SVD [AGM12]. Switching [FPRC13, Pot10]. Sylvester [SS10b]. Symmetric [BWY12]. Symmetry [BEPS12]. Symposium [IEE10, IEE11, IEE12, IEE13]. Systems [AE11, AY10, Cha12, KMP10, KMP11, LS13].

Table [Ano10i, Ano11j, Ano12j, Ano13n]. Tabulation [Tho13]. Tardos [HS13]. Tasks [ABGG12]. Technique [BP12]. Techniques [CEGP10, Dwo11]. Temperature [DPR^{+}10]. Test [HM10, LR12, SS10b]. Testable [BSRZS12, BWY12, HL13]. Testers [Soh12]. Testing [BBBY12, BKS^{+}10, BGS10, CMOS11, FS13, HSS11, JR11, Mul12, TR10]. Theorem [CC10, FF11, IKM10, JPY12, Kan12, iKR10, CGS12]. Theorems [CD11, EJT10, TW11, WN11]. Theoretic [CM11]. Theory [Mul12]. Thorup [PR10]. Three [GKKS13, Vid13b, NNN12]. Three-Player [Vid13b]. Threshold [COV13, DKN10, IPS13, Kan11b, Kan12, Sly10]. Tight [AE11, BDSS11, BEO^{+}13, BFNS12, CLN13, FW12, SS12a]. Tile [DLP^{+}12]. Time [BG13, BGS11, BSWN10, BKM^{+}11, BFNS12, CMT11, CLN13, CNP^{+}11, DP10, FS13, KS11a, KMP11, MS13, PSSZ10, SS10a, She13, Val12]. Title [Ano10j, Ano10k, Ano11k, Ano11l, Ano12k, Ano12l, Ano13o, Ano13p]. Tolerance [DPR^{+}10]. Tools [KW12]. Topic [AGM12]. Topological [AE11]. Topologies [SS12a]. Tractable [CLX10, CMPP13, iKT11, KT11]. Tradeoffs [CCKM10]. Transform [BBDS12]. Transition [Sly10]. Traveling [GSS11, Vis12]. Tree [GP10, KSS13, NPS11]. Trees [BIL12, BDR^{+}12, ES11]. Treewidth [BDD^{+}13, CNP^{+}11, GS13b]. Triangle [WW10]. Triangles [RW11]. Triangulations [Rub13]. Truthful [HK12, KSM10]. TSP [BG13, MS11a, MS11b]. Tutorial [Dwo11]. Two [IPS13, JPY12, KS11b, Li12, MMP^{+}10, SS12b]. Two-Party [JPY12, MMP^{+}10]. Two-Source [Li12]. Two-Spin [SS12b].

Uncertainty [BGKS13]. Unconstrained [BFNS12]. Understanding [CDW13]. Undirected [BFM13, Bjö10, GS13a, Mad10]. Unified [BGS10, FNS11]. uniform [CHK12]. Unique [ABS10a, BP13, Her11, KMM11]. Unique-SAT [Her11]. Uniqueness [Sly10]. Unit [Rub13]. Universal [BDR^{+}12, DLP^{+}12]. Universally [BN10]. Unprovability [CLP13a]. Unsplittable [BSW11]. Update [BGS11]. Upper [BK10]. USA [IEE10, IEE11, IEE12, IEE13]. Using [CCH^{+}12, GH11, SST11]. Utility [LD11].

V [Mul12]. Value [CD11]. Valued [TZ12]. Variables [DDO^{+}13]. Vector [DGY10]. Vegas [IEE10]. Versions [EJT10]. Vertex [CLLM10, CMPP13, KSM10, LRV13, MM10]. Vertices [KW12]. Via [DPV11, AKO11, And10, BRS11, CVZ10, CEGP10, CV13a, CDK12, Cyg13, DG13, DLWZ11, GW12, HLP13, KLL^{+}12, Kop12, NN13, PTW10, Wee10, WY10]. Volume [DR10a]. Voronoi [CHLX13].

Walking [LM12]. Walks [CMOS11]. Way [COPV13, HO11, iKT11, KT11]. Weight [AW13, DP10, RT12, RW11]. Weight-Scaling [RT12]. Weighted [CV13b, HLP13, KSS13, NPS11]. Weights [CC10, HR10]. Weil [KL11]. Welfare [BGMS11, CDW12a, HK12]. Which [BEK^{+}11]. Windows [BCM13]. within [Rot13]. without [GH11, PSW13]. Work [BG11]. Worlds [BGKS13]. Worst [ANS10]. Worst-Case [ANS10].

XOR [Vid13b]. XOS [FV13].

Yields [ABS10b].

Zero [CLP13b, KLL^{+}12]. Zero-Communication [KLL^{+}12]. Zwick [PR10].

References

Alistarh:2011:CR

[AAGG11] Dan Alistarh, James Aspnes, Seth Gilbert, and Rachid Guerraoui. The complexity of renaming. In IEEE [IEE11], pages 718–727. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Aharonov:2011:ALC

[AALV11] Dorit Aharonov, Itai Arad, Zeph Landau, and Umesh Vazirani. The 1D area law and the complexity of quantum states: a combinatorial approach. In IEEE [IEE11], pages 324–333. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Andrews:2010:MCN

[AAZ10] Matthew Andrews, Spyridon Antonakopoulos, and Lisa Zhang. Minimum-cost network design with (dis)economies of scale. In IEEE [IEE10], pages 585–592. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Alistarh:2012:HAT

[ABGG12] D. Alistarh, M. A. Bender, S. Gilbert, and R. Guerraoui. How to allocate tasks asynchronously. In IEEE [IEE12], pages 331–340. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375311.

IEEE Computer Society order number P????.

Arora:2010:SAU

[ABS10a] Sanjeev Arora, Boaz Barak, and David Steurer. Subexponential algorithms for unique games and related problems. In IEEE [IEE10], pages 563–572. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Awasthi:2010:SYP

[ABS10b] Pranjal Awasthi, Avrim Blum, and Or Sheffet. Stability yields a PTAS for k-median and k-means clustering. In IEEE [IEE10], pages 309–318. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Andoni:2011:NLL

[ACNN11] Alexandr Andoni, Moses S.

Charikar, Ofer Neiman, and Huy L. Nguyen. Near linear lower bound for dimension reduction in L1. In IEEE [IEE11], pages 315–323. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Aharonov:2011:CCL

[AE11] Dorit Aharonov and Lior Eldar. On the complexity of commuting local hamiltonians, and tight conditions for topological order in such systems. In IEEE [IEE11], pages 334–343. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Avigdor-Elgrabli:2013:ICA

[AER13] N. Avigdor-Elgrabli and Y. Rabani. An improved competitive algorithm for reordering buffer management. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Alaei:2013:SEA

[AFHH13] S. Alaei, Hu Fu, N. Haghpanah, and J. Hartline. The simple economics of approximately optimal auctions. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Arora:2012:LTM

[AGM12] S. Arora, Rong Ge, and A. Moitra. Learning topic models – going beyond SVD. In IEEE [IEE12], pages 1– 10. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375276.

IEEE Computer Society order number P????.

Arora:2013:TBA

[AGS13] S. Arora, Rong Ge, and A. K.

Sinop. Towards a better approximation for sparsest cut? In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Applebaum:2011:HGA

[AIK11] Benny Applebaum, Yuval Ishai, and Eyal Kushilevitz. How to garble arithmetic circuits. In IEEE [IEE11], pages 120–129. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Andoni:2010:PAE

[AKO10] Alexandr Andoni, Robert Krauthgamer, and Krzysztof Onak. Polylogarithmic approximation for edit distance and the asymmetric query complexity. In IEEE [IEE10], pages 377–386. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Andoni:2011:SAP

[AKO11] Alexandr Andoni, Robert Krauthgamer, and Krzysztof Onak. Streaming algorithms via precision sampling. In IEEE [IEE11], pages 363–372. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Alaei:2011:BCA

[Ala11] Saeed Alaei. Bayesian combinatorial auctions: Expanding single buyer mechanisms to many buyers. In IEEE [IEE11], pages 512–521. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Alon:2010:NLL

[Alo10] Noga Alon. A non-linear lower bound for planar epsilon-nets. In IEEE [IEE10], pages 341–346. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Agarwal:2012:RSS

[AMS12] P. K. Agarwal, J. Matouek, and M. Sharir. On range searching with semialgebraic sets II. In IEEE [IEE12], pages 420–429. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375320.

IEEE Computer Society order number P????.

Andrews:2010:AAE

[And10] Matthew Andrews. Approximation algorithms for the edge-disjoint paths problem via Raecke decompositions. In IEEE [IEE10], pages 277–286. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:AI

[Ano10a] Anonymous. Author index. In IEEE [IEE10], pages 824–826. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:CP

[Ano10b] Anonymous. Copyright page. In IEEE [IEE10], page iv. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:CA

[Ano10c] Anonymous. Cover art. In IEEE [IEE10], page C1. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:F

[Ano10d] Anonymous. Foreword. In IEEE [IEE10], page xii. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:OC

[Ano10e] Anonymous. Organizing committee. In IEEE [IEE10], page xiii. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:PC

[Ano10f] Anonymous. Program committee. In IEEE [IEE10], page xiv. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:R

[Ano10g] Anonymous. Reviewers. In IEEE [IEE10], pages xv–xvi. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:RP

[Ano10h] Anonymous. Roster page.

In IEEE [IEE10], page 828. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:TC

[Ano10i] Anonymous. Table of contents. In IEEE [IEE10], pages v– x. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:TP

[Ano10j] Anonymous. Title page i. In IEEE [IEE10], page i. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2010:TPI

[Ano10k] Anonymous. Title page iii. In IEEE [IEE10], page iii. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Anonymous:2011:AI

[Ano11a] Anonymous. Author index. In IEEE [IEE11], pages 847–849. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.

org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:A

[Ano11b] Anonymous.  Awards.  In

IEEE [IEE11], page xvii. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:CP

[Ano11c] Anonymous. Copyright page.

In IEEE [IEE11], page iv. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:CA

[Ano11d] Anonymous. Cover art. In IEEE [IEE11], page C1. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:F

[Ano11e] Anonymous. Foreword. In IEEE [IEE11], page xii. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:OC

[Ano11f] Anonymous. Organizing committee. In IEEE [IEE11], page xiii. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:PC

[Ano11g] Anonymous. Program committee. In IEEE [IEE11], page xiv. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:PI

[Ano11h] Anonymous. Publisher’s information. In IEEE [IEE11], page 850. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:R

[Ano11i] Anonymous. Reviewers. In IEEE [IEE11], pages xv–xvi. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:TC

[Ano11j] Anonymous. Table of contents. In IEEE [IEE11], pages v– xi. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:TP

[Ano11k] Anonymous. Title page i. In IEEE [IEE11], page i. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2011:TPI

[Ano11l] Anonymous. Title page iii.

In IEEE [IEE11], page iii. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Anonymous:2012:AI

[Ano12a]  Anonymous.  Author index.

In IEEE [IEE12], pages 768– 770. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375356.

IEEE Computer Society order number P????.

Anonymous:2012:A

[Ano12b] Anonymous. Awards. In IEEE [IEE12], page xix. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375365.

IEEE Computer Society order number P????.

Anonymous:2012:BC

[Ano12c]  Anonymous.  Back cover.

In IEEE [IEE12], page C4. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375366.

IEEE Computer Society order number P????.

Anonymous:2012:CN

[Ano12d] Anonymous. Copyright notice. In IEEE [IEE12], page iv. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375360.

IEEE Computer Society order number P????.

Anonymous:2012:F

[Ano12e] Anonymous. Foreword. In IEEE [IEE12], pages xii–xiii. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375361.

IEEE Computer Society order number P????.

Anonymous:2012:OC

[Ano12f] Anonymous. Organizing committee. In IEEE [IEE12], page xiv. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375362.

IEEE Computer Society order number P????.

Anonymous:2012:PC

[Ano12g] Anonymous. Program committee. In IEEE [IEE12], page xv. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375363.

IEEE Computer Society order number P????.

Anonymous:2012:PI

[Ano12h] Anonymous. [publisher’s information]. In IEEE [IEE12], page 772. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375357.

IEEE Computer Society order number P????.

Anonymous:2012:R

[Ano12i] Anonymous. Reviewers. In IEEE [IEE12], pages xvi– xviii. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375364.

IEEE Computer Society order number P????.

Anonymous:2012:TC

[Ano12j] Anonymous. Table of contents. In IEEE [IEE12], pages v–xi. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375367.

IEEE Computer Society order number P????.

Anonymous:2012:TPa

[Ano12k] Anonymous. Title page. In IEEE [IEE12], page i. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375358.

IEEE Computer Society order number P????.

Anonymous:2012:TPb

[Ano12l] Anonymous. Title page. In IEEE [IEE12], page iii. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/

stamp.jsp?arnumber=6375359.

IEEE Computer Society order number P????.

Anonymous:2013:AI

[Ano13a] Anonymous. Author index. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:A

[Ano13b] Anonymous. Awards. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:CP

[Ano13c] Anonymous. Copyright page. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:CA

[Ano13d] Anonymous. Cover art. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.

                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:F

[Ano13e] Anonymous. Foreword. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:OCS

[Ano13f] Anonymous. Organizing committee and sponsors. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:PAa

[Ano13g] Anonymous. Proceedings available. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:PAb

[Ano13h] Anonymous. Proceedings available. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428.

LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:PC

[Ano13i] Anonymous. Program committee. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:PI

[Ano13j] Anonymous. Publisher’s information. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:QLa

[Ano13k] Anonymous. Quick links. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:QLb

[Ano13l] Anonymous. Quick links. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:R

[Ano13m] Anonymous. Reviewers. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:TC

[Ano13n] Anonymous. Table of contents. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:TP

[Ano13o] Anonymous. Title page i. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Anonymous:2013:TPI

[Ano13p] Anonymous. Title page iii. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Arbitman:2010:BCH

[ANS10] Yuriy Arbitman, Moni Naor, and Gil Segev. Backyard cuckoo hashing: Constant worstcase operations with a succinct representation. In IEEE [IEE10], pages 787–796. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Adamaszek:2013:ASM

[AW13] A. Adamaszek and A. Wiese.

Approximation schemes for maximum weight independent set of rectangles. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Alon:2010:SLS

[AY10] Noga Alon and Raphael Yuster.

Solving linear systems through nested dissection. In IEEE [IEE10], pages 225–234. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Ben-Aroya:2010:LLD

[BAETS10] Avraham Ben-Aroya, Klim Efremenko, and Amnon Ta-Shma. Local list decoding with a constant number of queries. In IEEE [IEE10], pages 715–722. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Bansal:2010:CAD

[Ban10] Nikhil Bansal. Constructive algorithms for discrepancy minimization. In IEEE [IEE10], pages 3–10. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Balcan:2012:APT

[BBBY12] M.-F. Balcan, E. Blais, A. Blum, and Liu Yang. Active property testing. In IEEE [IEE12], pages 21–30. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375278.

IEEE Computer Society order number P????.

Blocki:2012:JLT

[BBDS12] J. Blocki, A. Blum, A. Datta, and O. Sheffet. The Johnson– Lindenstrauss Transform itself preserves differential privacy. In IEEE [IEE12], pages 410–419. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375319.

IEEE Computer Society order number P????.

Bansal:2011:PCA

[BBMN11] Nikhil Bansal, Niv Buchbinder, Aleksander Madry, and Joseph (Seffi) Naor. A polylogarithmic-competitive algorithm for the k-server problem. In IEEE [IEE11], pages 267–276. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Benoit:2012:QOM

[BBvdH12] A. Benoit, A. Bostan, and J. van der Hoeven. Quasioptimal multiplication of linear differential operators. In IEEE [IEE12], pages 524– 530. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375331.

IEEE Computer Society order number P????.

Beame:2013:EDF

[BCM13] P. Beame, R. Clifford, and W. Machmouchi. Element distinctness, frequency moments, and sliding windows. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Babai:2013:FCF

[BCS^{+}13] L. Babai, Xi Chen, Xiaorui Sun, Shang-Hua Teng, and J. Wilmes. Faster canonical forms for strongly regular graphs. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Bodlaender:2013:CAA

[BDD^{+}13] H. L. Bodlaender, P. G. Drange, M. S. Dregi, F. V. Fomin, D. Lokshtanov, and M. Pilipczuk. An O(c^{k}n) 5approximation algorithm for treewidth. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Boutsidis:2011:NOC

[BDMI11] Christos Boutsidis, Petros Drineas, and Malik Magdon-Ismail. Near optimal column-based matrix reconstruction. In IEEE [IEE11], pages 305–314. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Busch:2012:SJS

[BDR^{+}12] C. Busch, C. Dutta, J. Radhakrishnan, R. Rajaraman, and S. Srinivasagopalan. Split and join: Strong partitions and universal Steiner trees for graphs. In IEEE [IEE12], pages 81–90. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375285.

IEEE Computer Society order number P????.

Bhattacharyya:2011:TLB

[BDSS11] Arnab Bhattacharyya, Zeev Dvir, Amir Shpilka, and Shubhangi Saraf. Tight lower bounds for 2-query LCCs over finite fields. In IEEE [IEE11], pages 638–647. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Blume:2011:WNL

[BEK^{+}11] Lawrence Blume, David Easley, Jon Kleinberg, Robert Kleinberg, and Eva Tardos. Which networks are least susceptible to cascading failures? In IEEE [IEE11], pages 393–402. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Belovs:2012:LGB

[Bel12]  A. Belovs.  Learning-graph-

based quantum algorithm for kdistinctness. In IEEE [IEE12], pages 207–216. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375298.

IEEE Computer Society order number P????.

Braverman:2013:TBS

[BEO^{+}13] M. Braverman, F. Ellen, R. Oshman, T. Pitassi, and V. Vaikuntanathan. A tight bound for set disjointness in the messagepassing model. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Barenboim:2012:LDS

[BEPS12] L. Barenboim, M. Elkin, S. Pettie, and J. Schneider. The locality of distributed symmetry breaking. In IEEE [IEE12], pages 321–330. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375310.

IEEE Computer Society order number P????.

Berkholz:2012:CFN

[Ber12] C. Berkholz. On the complexity of finding narrow proofs. In IEEE [IEE12], pages 351– 360. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375313.

IEEE Computer Society order number P????.

Bansal:2011:MMG

[BFK^{+}11] Nikhil Bansal, Uriel Feige, Robert Krauthgamer, Konstantin Makarychev, Viswanath Nagarajan, Joseph (Seffi) Naor, and Roy Schwartz. Minmax graph partitioning and small set expansion. In IEEE [IEE11], pages 17–26. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Bilo:2013:PSU

[BFM13] V. Bilo, M. Flammini, and L. Moscardelli. The price of stability for undirected broadcast network design with fair cost allocation is constant. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Buchbinder:2012:TLT

[BFNS12] N. Buchbinder, M. Feldman, J. Naor, and R. Schwartz. A tight linear time (1/2 )- approximation for unconstrained submodular maximization. In IEEE [IEE12], pages 649– 658. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375344.

IEEE Computer Society order number P????.

Braun:2012:ALL

[BFPS12] G. Braun, S. Fiorini, S. Pokutta, and D. Steurer. Approximation limits of linear programs (beyond hierarchies). In IEEE [IEE12], pages 480– 489. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375326.

IEEE Computer Society order number P????.

Bender:2011:MEL

[BG11] Michael A. Bender and Seth Gilbert. Mutual exclusion with O(log^{2} log n) amortized work. In IEEE [IEE11], pages 728–737. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Bartal:2013:LTA

[BG13] Y. Bartal and L.-A. Gottlieb.

A linear time approximation scheme for Euclidean TSP. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Barak:2012:MLC

[BGH^{+}12] B. Barak, P. Gopalan, J. Hastad, R. Meka, P. Raghavendra, and D. Steurer. Making the long code shorter. In IEEE [IEE12], pages 370–379. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375315.

IEEE Computer Society order number P????.

Bassily:2013:CWP

[BGKS13] R. Bassily, A. Groce, J. Katz, and A. Smith. Coupledworlds privacy: Exploiting adversarial uncertainty in statistical data privacy. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Blum:2011:WPM

[BGMS11] Avrim Blum, Anupam Gupta, Yishay Mansour, and Ankit Sharma. Welfare and profit maximization with production costs. In IEEE [IEE11], pages 77– 86. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Bhattacharyya:2010:UFT

[BGS10] Arnab Bhattacharyya, Elena Grigorescu, and Asaf Shapira. A unified framework for testing linear-invariant properties. In IEEE [IEE10], pages 478–487. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Baswana:2011:FDM

[BGS11] Surender Baswana, Manoj Gupta, and Sandeep Sen. Fully dynamic maximal matching in O(log n) update time. In IEEE [IEE11], pages 383–392. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Bjorklund:2013:PDH

[BH13] A. Bjorklund and T. Husfeldt. The parity of directed Hamiltonian cycles. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Beck:2012:LDB

[BIL12] C. Beck, R. Impagliazzo, and S. Lovett. Large deviation bounds for decision trees and sampling lower bounds for AC0circuits. In IEEE [IEE12], pages 101–110. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375287.

IEEE Computer Society order number P????.

Bjorklund:2010:DSU

[Bjö10] Andreas Björklund. Determinant sums for undirected Hamiltonicity. In IEEE [IEE10], pages 173–182. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Bosek:2010:SEU

[BK10] Bartlomiej Bosek and Tomasz Krawczyk. The sub-exponential upper bound for on-line chain partitioning. In IEEE [IEE10], pages 347–354. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Brakerski:2012:EIC

[BK12] Z. Brakerski and Y. T. Kalai. Efficient interactive coding against adversarial noise. In IEEE [IEE12], pages 160– 166. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375293.

IEEE Computer Society order number P????.

Brakerski:2010:OHB

[BKKV10] Zvika Brakerski, Yael Tauman Kalai, Jonathan Katz, and Vinod Vaikuntanathan. Overcoming the hole in the bucket: Public-key cryptography resilient to continual memory leakage. In IEEE [IEE10], pages 501–510. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Blasiak:2011:LPP

[BKL11] Anna Blasiak, Robert Kleinberg, and Eyal Lubetzky. Lexicographic products and the power of non-linear network coding. In IEEE [IEE11], pages 609–618. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Borradaile:2011:MSM

[BKM^{+}11] Glencora Borradaile, Philip N.

Klein, Shay Mozes, Yahav Nussbaum, and Christian Wulff-Nilsen. Multiple-source multiplesink maximum flow in directed planar graphs in nearlinear time. In IEEE [IEE11], pages 170–179. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Bindel:2011:HBF

[BKO11] David Bindel, Jon Kleinberg, and Sigal Oren. How bad is forming your own opinion? In IEEE [IEE11], pages 57– 66. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Bhattacharyya:2010:OTR

[BKS^{+}10] Arnab Bhattacharyya, Swastik Kopparty, Grant Schoenebeck, Madhu Sudan, and David Zuckerman. Optimal testing of Reed–Muller codes. In IEEE [IEE10], pages 488–497. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Badanidiyuru:2013:BK

[BKS13] A. Badanidiyuru, R. Kleinberg, and A. Slivkins. Bandits with knapsacks. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Braverman:2011:GCS

[BMMN11] Mark Braverman, Konstantin Makarychev, Yury Makarychev, and Assaf Naor. The Grothendieck constant is strictly smaller than Krivine’s bound. In IEEE [IEE11], pages 453–462. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Brenner:2010:IDP

[BN10] Hai Brenner and Kobbi Nissim. Impossibility of differentially private universally optimal mechanisms. In IEEE [IEE10], pages 71–80. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Bansal:2010:GS

[BP10] Nikhil Bansal and Kirk Pruhs.

The geometry of scheduling. In IEEE [IEE10], pages 407–414.

ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Bitansky:2012:ION

[BP12] N. Bitansky and O. Paneth. From the impossibility of obfuscation to a new non-black-box simulation technique. In IEEE [IEE12], pages 223– 232. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375300.

IEEE Computer Society order number P????.

Braun:2013:CIU

[BP13] G. Braun and S. Pokutta. Common information and unique disjointness. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Bogdanov:2011:PRO

[BPW11] Andrej Bogdanov, Periklis A. Papakonstaninou, and Andrew Wan. Pseudorandomness for read-once formulas. In IEEE [IEE11], pages 240–246. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Braverman:2011:IEA

[BR11] Mark Braverman and Anup Rao. Information equals amortized communication. In IEEE [IEE11], pages 748–757. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Braverman:2010:PGR

[BRRY10] Mark Braverman, Anup Rao, Ran Raz, and Amir Yehudayoff. Pseudorandom generators for regular branching programs. In IEEE [IEE10], pages 40– 47. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Barak:2011:RSP

[BRS11] Boaz Barak, Prasad Raghavendra, and David Steurer. Rounding semidefinite programming hierarchies via global correlation. In IEEE [IEE11], pages 472–481. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Braverman:2013:DPC

[BRWY13] M. Braverman, A. Rao, O. Weinstein, and A. Yehudayoff. Direct products in communication complexity. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Belkin:2010:PLD

[BS10] Mikhail Belkin and Kaushik Sinha. Polynomial learning of distribution families. In IEEE [IEE10], pages 103–112. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Ben-Sasson:2013:CRP

[BSKK^{+}13] E. Ben-Sasson, Y. Kaplan, S. Kopparty, O. Meir, and H. Stichtenoth. Constant rate PCPs for circuit-SAT with sublinear query complexity. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428.

LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Ben-Sasson:2012:ACA

[BSLRZ12] E. Ben-Sasson, S. Lovett, and N. Ron-Zewi. An additive combinatorics approach relating rank to communication complexity. In IEEE [IEE12], pages 177–186. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375295.

IEEE Computer Society order number P????.

Ben-Sasson:2012:SAI

[BSRZS12] E. Ben-Sasson, N. Ron-Zewi, and M. Sudan. Sparse affineinvariant linear codes are locally testable. In IEEE [IEE12], pages 561–570. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375335.

IEEE Computer Society order number P????.

Bonsma:2011:CFA

[BSW11] Paul Bonsma, Jens Schulz, and Andreas Wiese. A constant factor approximation algorithm for unsplittable flow on paths. In IEEE [IEE11], pages 47– 56. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.

                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Borradaile:2010:MSC

[BSWN10] Glencora Borradaile, Piotr Sankowski, and Christian Wulff-Nilsen. Min st-cut oracle for planar graphs with nearlinear preprocessing time. In IEEE [IEE10], pages 601–610. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Brody:2010:CPP

[BV10] Joshua Brody and Elad Verbin. The coin problem and pseudorandomness for branching programs. In IEEE [IEE10], pages 30–39. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Brakerski:2011:EFH

[BV11] Zvika Brakerski and Vinod Vaikuntanathan. Efficient fully homomorphic encryption from (standard) LWE. In IEEE [IEE11], pages 97–106. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Blais:2012:PSF

[BWY12] E. Blais, A. Weinstein, and Y. Yoshida. Partially symmetric functions are efficiently isomorphism-testable. In IEEE [IEE12], pages 551–560. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=

6375334. IEEE Computer Society order number P????.

Cai:2010:DDT

[CC10] Jin-Yi Cai and Xi Chen. A decidable dichotomy theorem on directed graph homomorphisms with non-negative weights. In IEEE [IEE10], pages 437–446. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Chitnis:2012:DFA

[CCH^{+}12] R. Chitnis, M. Cygan, M. Hajiaghayi, M. Pilipczuk, and M. Pilipczuk. Designing FPT algorithms for cut problems using randomized contractions. In IEEE [IEE12], pages 460– 469. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375324.

IEEE Computer Society order number P????.

Chakrabarti:2010:ICT

[CCKM10] Amit Chakrabarti, Graham Cormode, Ranganath Kondapally, and Andrew McGregor. Information cost tradeoffs for augmented index and streaming language recognition. In IEEE [IEE10], pages 387–396. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Cai:2011:EVT

[CD11] Yang Cai and Constantinos Daskalakis. Extreme-value theorems for optimal multidimensional pricing. In IEEE [IEE11], pages 522–531. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chlamtac:2012:ESS

[CDK12] E. Chlamtać, M. Dinitz, and R. Krauthgamer. Everywheresparse spanners via dense subgraphs. In IEEE [IEE12], pages 758–767. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375355.

IEEE Computer Society order number P????.

Cai:2012:OMD

[CDW12a] Yang Cai, C. Daskalakis, and S. M. Weinberg. Optimal multidimensional mechanism design: Reducing revenue to welfare maximization. In IEEE [IEE12], pages 130–139. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375290.

IEEE Computer Society order number P????.

Christandl:2012:CML

[CDW12b] M. Christandl, B. Doran, and M. Walter. Computing multiplicities of Lie group representations. In IEEE [IEE12], pages 639–648. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375343.

IEEE Computer Society order number P????.

Cai:2013:UIM

[CDW13] Yang Cai, C. Daskalakis, and S. M. Weinberg. Understanding incentives: Mechanism design becomes algorithm design. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.

                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Chekuri:2011:AAS

[CE11] Chandra Chekuri and Alina Ene. Approximation algorithms for submodular multiway partition. In IEEE [IEE11], pages 807–816. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chen:2010:FMD

[CEGP10] Ning Chen, Edith Elkind, Nick Gravin, and Fedor Petrov. Frugal mechanism design via spectral techniques. In IEEE [IEE10], pages 755–764. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Caragiannis:2011:ECA

[CFGS11] Ioannis Caragiannis, Angelo Fanelli, Nick Gravin, and Alexander Skopalik. Efficient computation of approximate pure Nash equilibria in congestion games. In IEEE [IEE11], pages 532–541. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Cygan:2012:AAB

[CGS12] M. Cygan, H. N. Gabow, and P. Sankowski. Algorithmic applications of Baur–Strassen’s Theorem: Shortest cycles, diameter and matchings. In IEEE [IEE12], pages 531–540. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375332.

IEEE Computer Society order number P????.

Chazelle:2012:DIS

[Cha12]  B. Chazelle.  The dynam-

ics of influence systems. In IEEE [IEE12], pages 311– 320. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375309.

IEEE Computer Society order number P????.

Chan:2013:KMP

[Cha13] T. M. Chan. Klee’s measure problem made easy. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Cygan:2012:LRC

[CHK12] M. Cygan, M. Hajiaghayi, and S. Khuller. LP rounding for kcenters with non-uniform hard capacities. In IEEE [IEE12], pages 273–282. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375305.

IEEE Computer Society order number P????.

Chen:2013:CIV

[CHLX13] D. Z. Chen, Ziyun Huang, Yangwei Liu, and Jinhui Xu. On clustering induced Voronoi diagrams. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Clarkson:2010:SOM

[CHW10] Kenneth L. Clarkson, Elad Hazan, and David P. Woodruff. Sublinear optimization for machine learning. In IEEE [IEE10], pages 449–457. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Chailloux:2011:OBQ

[CK11] Andre Chailloux and Iordanis Kerenidis. Optimal bounds for quantum bit commitment. In IEEE [IEE11], pages 354–362. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chierichetti:2011:ML

[CKR11] Flavio Chierichetti, Ravi Kumar, and Prabhakar Raghavan. Markov layout. In IEEE [IEE11], pages 492–501. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chuzhoy:2012:PAA

[CL12] J. Chuzhoy and Shi Li. A polylogarithmic approximation algorithm for edge-disjoint paths with congestion 2. In IEEE [IEE12], pages 233–242. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=

6375301. IEEE Computer Society order number P????.

Cheung:2011:GCN

[CLL11] Ho Yee Cheung, Lap Chi Lau, and Kai Man Leung. Graph connectivities, network coding, and expander graphs. In IEEE [IEE11], pages 190–199. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Charikar:2010:VSA

[CLLM10] Moses Charikar, Tom Leighton, Shi Li, and Ankur Moitra. Vertex sparsifiers and abstract rounding algorithms. In IEEE [IEE10], pages 265–274. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Chalermsook:2013:ISI

[CLN13] P. Chalermsook, B. Laekhanukit, and D. Nanongkai. Independent set, induced matching, and pricing: Connections and tight (subexponential time) approximation hardnesses. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Canetti:2010:AHC

[CLP10] Ran Canetti, Huijia Lin, and Rafael Pass. Adaptive hardness and composable security in the plain model from standard assumptions. In IEEE [IEE10], pages 541–550. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Canetti:2013:UEF

[CLP13a] R. Canetti, Huijia Lin, and R. Pass. From unprovability to environmentally friendly protocols. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Chung:2013:CRC

[CLP13b] Kai-Min Chung, Huijia Lin, and R. Pass. Constantround concurrent zero knowledge from P -certificates. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Chan:2013:ACS

[CLRS13] Siu On Chan, J. R. Lee, P. Raghavendra, and D. Steurer. Approximate constraint satisfaction requires large LP relaxations. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Cai:2010:HAM

[CLX10] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holographic algorithms with matchgates capture precisely tractable planar_{#}CSP. In IEEE [IEE10], pages 427–436. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Chen:2011:MDS

[CM11] Jing Chen and Silvio Micali. Mechanism design with settheoretic beliefs. In IEEE [IEE11], pages 87–96. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Czumaj:2011:PGR

[CMOS11] Artur Czumaj, Morteza Monemizadeh, Krzysztof Onak, and Christian Sohler. Planar graphs: Random walks and bipartiteness testing. In IEEE [IEE11], pages 423–432. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Cygan:2013:PDK

[CMPP13] M. Cygan, D. Marx, M. Pilipczuk, and M. Pilipczuk. The planar directed K-vertex-disjoint paths problem is fixed-parameter tractable. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Caputo:2011:SMT

[CMT11] Pietro Caputo, Fabio Martinelli, and Fabio Lucio Toninelli. Sharp mixing time bounds for sampling random surfaces. In IEEE [IEE11], pages 130–139. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Cygan:2011:SCP

[CNP^{+}11] Marek Cygan, Jesper Nederlof, Marcin Pilipczuk, Michal Pilipczuk, Joham M. M. van Rooij, and Jakub Onufry Wojtaszczyk. Solving connectivity problems parameterized by treewidth in single exponential time. In IEEE [IEE11], pages 150–159. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chung:2013:SRO

[COPV13] Kai-Min Chung, R. Ostrovsky, R. Pass, and I. Visconti. Simultaneous resettability from one-way functions. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Coja-Oghlan:2013:CKC

[COV13] A. Coja-Oghlan and D. Vilenchik. Chasing the Kcolorability threshold. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/

FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Chung:2011:RCP

[CP11] Kai-Min Chung and Rafael Pass. The randomness complexity of parallel repetition. In IEEE [IEE11], pages 658–667. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chung:2013:KPI

[CPT13] Kai-Min Chung, R. Pass, and S. Telang. Knowledge-preserving interactive coding. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Celis:2011:BBS

[CRSW11] L. Elisa Celis, Omer Reingold, Gil Segev, and Udi Wieder. Balls and bins: Smaller hash families and faster evaluation. In IEEE [IEE11], pages 599–608. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Chattopadhyay:2012:LBI

[CS12] A. Chattopadhyay and R. Santhanam. Lower bounds on interactive compressibility by constant-depth circuits. In IEEE [IEE12], pages 619– 628. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375341.

IEEE Computer Society order number P????.

Chekuri:2013:AAE

[CS13] C. Chekuri and A. Sidiropoulos. Approximation algorithms for Euler genus and related problems. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Cheriyan:2013:AMC

[CV13a] J. Cheriyan and L. A. Vegh. Approximating minimum-cost knode connected subgraphs via independence-free graphs. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Chiplunkar:2013:RMA

[CV13b] A. Chiplunkar and S. Vishwanathan. On randomized memoryless algorithms for the weighted K-server problem. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Chandrasekaran:2012:CPM

[CVV12] K. Chandrasekaran, L. A. Vegh, and S. Vempala. The cutting plane method is polynomial for perfect matchings. In IEEE [IEE12], pages 571– 580. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375336.

IEEE Computer Society order number P????.

Chekuri:2010:DRR

[CVZ10] Chandra Chekuri, Jan Vondrak, and Rico Zenklusen. Dependent randomized rounding via exchange properties of combinatorial structures. In IEEE [IEE10], pages 575–584. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Cygan:2013:IAD

[Cyg13]  M. Cygan.  Improved ap-

proximation for 3-dimensional matching via bounded pathwidth local search. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Daskalakis:2013:LSI

[DDO^{+}13] C. Daskalakis, I. Diakonikolas, R. ODonnell, R. A. Servedio, and Li-Yang Tan. Learning sums of independent integer random variables. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

DiBattista:2010:QNP

[DFP10] Giuseppe Di Battista, Fabrizio Frati, and Janos Pach. On the queue number of planar graphs. In IEEE [IEE10], pages 365–374. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dinur:2013:PLD

[DG13] I. Dinur and V. Guruswami.

PCPs via low-degree long code and hardness for constrained hypergraph coloring. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Dvir:2010:MVC

[DGY10] Zeev Dvir, Parikshit Gopalan, and Sergey Yekhanin. Matching vector codes. In IEEE [IEE10], pages 705–714. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dodis:2010:CAC

[DHLAW10] Yevgeniy Dodis, Kristiyan Haralambiev, Adriana Lopez-Alt, and Daniel Wichs. Cryptography against continuous memory attacks. In IEEE [IEE10], pages 511–520. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Duchi:2013:LPS

[DJW13] J. C. Duchi, M. I. Jordan, and M. J. Wainwright. Local privacy and statistical minimax rates. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Diakonikolas:2010:BIF

[DKN10] Ilias Diakonikolas, Daniel M. Kane, and Jelani Nelson. Bounded independence fools degree-2 threshold functions. In IEEE [IEE10], pages 11– 20. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dinur:2010:HFI

[DKPS10] Irit Dinur, Subhash Khot, Will Perkins, and Muli Safra. Hardness of finding independent sets in almost 3-colorable graphs. In IEEE [IEE10], pages 212–221. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dvorak:2010:DFO

[DKT10] Zdenek Dvořak, Daniel Krał, and Robin Thomas. Deciding first-order properties for sparse graphs. In IEEE [IEE10], pages 133–142. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Doty:2012:TAM

[DLP^{+}12] D. Doty, J. H. Lutz, M. J. Patitz, R. T. Schweller, S. M. Summers, and D. Woods. The tile assembly model is intrinsically universal. In IEEE [IEE12], pages 302–310. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375308.

IEEE Computer Society order number P????.

Dodis:2011:SSC

[DLWW11] Yevgeniy Dodis, Allison Lewko, Brent Waters, and Daniel Wichs. Storing secrets on continually leaky devices. In IEEE [IEE11], pages 688–697. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Dodis:2011:PAN

[DLWZ11] Yevgeniy Dodis, Xin Li, Trevor D. Wooley, and David Zuckerman. Privacy amplification and non-malleable extractors via character sums. In IEEE [IEE11], pages 668–677. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Dujmovic:2013:LSQ

[DMW13] V. Dujmovic, P. Morin, and D. R. Wood. Layered separators for queue layouts, 3D graph drawing and nonrepetitive coloring. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Dwork:2012:PAP

[DNV12] C. Dwork, M. Naor, and S. Vadhan. The privacy of the analyst and the power of the state. In IEEE [IEE12], pages 400–409. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375318.

IEEE Computer Society order number P????.

Duan:2010:AMW

[DP10] Ran Duan and Seth Pettie. Approximating maximum weight matching in near-linear time. In IEEE [IEE10], pages 673–682. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Doty:2010:SFT

[DPR^{+}10] David Doty, Matthew J. Patitz, Dustin Reishus, Robert T. Schweller, and Scott M. Summers. Strong fault-tolerance for self-assembly with fuzzy temperature. In IEEE [IEE10], pages 417–426. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dadush:2011:ELA

[DPV11] Daniel Dadush, Chris Peikert, and Santosh Vempala. Enumerative lattice algorithms in any norm via M-ellipsoid coverings. In IEEE [IEE11], pages 580–589. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Deshpande:2010:EVS

[DR10a] Amit Deshpande and Luis Rademacher. Efficient volume sampling for row/column subset selection. In IEEE [IEE10], pages 329–338. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dughmi:2010:BBR

[DR10b] Shaddin Dughmi and Tim Roughgarden. Black-box randomized reductions in algorithmic mechanism design. In IEEE [IEE10], pages 775–784. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Drucker:2012:NLC

[Dru12] A. Drucker. New limits to classical and quantum instance compression. In IEEE [IEE12], pages 609–618. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375340.

IEEE Computer Society order number P????.

Drucker:2013:NDP

[Dru13] A. Drucker. Nondeterministic direct product reductions and the success probability of SAT solvers. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Dwork:2010:BDP

[DRV10] Cynthia Dwork, Guy N. Rothblum, and Salil Vadhan. Boosting and differential privacy. In IEEE [IEE10], pages 51–60. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Dughmi:2011:LRM

[DV11] Shaddin Dughmi and Jan Vondrák. Limitations of randomized mechanisms for combinatorial auctions. In IEEE [IEE11], pages 502–511. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Dwork:2011:PDP

[Dwo11] Cynthia Dwork. The promise of differential privacy: a tutorial on algorithmic techniques. In IEEE [IEE11], pages 1–2. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Elberfeld:2010:LVT

[EJT10] Michael Elberfeld, Andreas Jakoby, and Till Tantau. Logspace versions of the theorems of Bodlaender and Courcelle. In IEEE [IEE10], pages 143–152. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Elkin:2011:SSL

[ES11] Michael Elkin and Shay Solomon. Steiner shallow-light trees are exponentially lighter than spanning ones. In IEEE [IEE11], pages 373–382. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Falik:2011:APR

[FF11] Dvir Falik and Ehud Friedgut.

An algebraic proof of a robust social choice impossibility theorem. In IEEE [IEE11], pages 413–422. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Fraigniaud:2011:LDD

[FKP11] Pierre Fraigniaud, Amos Korman, and David Peleg. Local distributed decision. In IEEE [IEE11], pages 708–717. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Fomin:2012:PDA

[FLMS12] F. V. Fomin, D. Lokshtanov, N. Misra, and S. Saurabh. Planar F -deletion: Approximation, kernelization and optimal FPT algorithms. In IEEE [IEE12], pages 470–479. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375325.

IEEE Computer Society order number P????.

Feldman:2011:UCG

[FNS11] Moran Feldman, Joseph (Seffi) Naor, and Roy Schwartz. A unified continuous greedy algorithm for submodular maximization. In IEEE [IEE11], pages 570–579. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Filmus:2013:ACL

[FPRC13] Y. Filmus, T. Pitassi, R. Robere, and S. A. Cook. Average case lower bounds for monotone switching networks. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Fukunaga:2012:IRA

[FR12] T. Fukunaga and R. Ravi. Iterative rounding approximation algorithms for degree-bounded node-connectivity network design. In IEEE [IEE12], pages 263–272. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375304.

IEEE Computer Society order number P????.

Forbes:2013:QTI

[FS13] M. A. Forbes and A. Shpilka.

Quasipolynomial-time identity testing of non-commutative and Read-once oblivious algebraic branching programs. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Feldman:2013:OBA

[FV13] V. Feldman and J. Vondrak.

Optimal bounds on approximation of submodular and XOS functions by juntas. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Filmus:2012:TCA

[FW12] Y. Filmus and J. Ward. A tight combinatorial algorithm for submodular maximization subject to a matroid constraint. In IEEE [IEE12], pages 659–668. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375345.

IEEE Computer Society order number P????.

Garg:2013:CIO

[GGH^{+}13] S. Garg, C. Gentry, S. Halevi, M. Raykova, A. Sahai, and B. Waters. Candidate indistinguishability obfuscation and functional encryption for all circuits. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gentry:2011:FHE

[GH11] Craig Gentry and Shai Halevi. Fully homomorphic encryption without squashing using depth-3 arithmetic circuits. In IEEE [IEE11], pages 107–109. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Garber:2013:PNL

[GH13] D. Garber and E. Hazan. Playing non-linear games with linear oracles. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Guruswami:2013:ESD

[GK13] V. Guruswami and S. Kopparty.

Explicit subspace designs. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gupta:2013:ACC

[GKKS13] A. Gupta, P. Kamath, N. Kayal, and R. Saptharishi. Arithmetic circuits: A chasm at depth three. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gupta:2011:ERR

[GKL11] Ankit Gupta, Neeraj Kayal, and Satya Lokam. Efficient reconstruction of random multilinear formulas. In IEEE [IEE11], pages 778–787. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Gopalan:2011:FKR

[GKM^{+}11] Parikshit Gopalan, Adam Klivans, Raghu Meka, Daniel Stefankovic, Santosh Vempala, and Eric Vigoda. An FP-TAS for #knapsack and related counting problems. In IEEE [IEE11], pages 817–826. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Garay:2013:RPD

[GKM^{+}13] J. Garay, J. Katz, U. Maurer, B. Tackmann, and V. Zikas. Rational protocol design: Cryptography against incentivedriven adversaries. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gupta:2011:AAC

[GKMR11] Anupam Gupta, Ravishankar Krishnaswamy, Marco Molinaro, and R. Ravi. Approximation algorithms for correlated knapsacks and nonmartingale bandits. In IEEE [IEE11], pages 827–836. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Goyal:2012:CNM

[GLOV12] V. Goyal, Chen-Kuei Lee, R. Ostrovsky, and I. Visconti. Constructing non-malleable commitments: A black-box approach. In IEEE [IEE12], pages 51–60. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375281.

IEEE Computer Society order number P????.

Gradwohl:2010:SRC

[GLR10] Ronen Gradwohl, Noam Livne, and Alon Rosen. Sequential rationality in cryptographic protocols. In IEEE [IEE10], pages 623–632. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Goyal:2011:SCP

[GM11] Vipul Goyal and Hemanta K.

Maji. Stateless cryptographic protocols. In IEEE [IEE11], pages 678–687. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Gopalan:2012:BPG

[GMR^{+}12] P. Gopalan, R. Meka, O. Reingold, L. Trevisan, and S. Vadhan. Better pseudorandom generators from milder pseudorandom restrictions. In IEEE [IEE12], pages 120– 129. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375289.

IEEE Computer Society order number P????.

Gelles:2011:EEC

[GMS11] Ran Gelles, Ankur Moitra, and Amit Sahai. Efficient and explicit coding for interactive communication. In IEEE [IEE11], pages 768–777. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Gosset:2013:QSQ

[GN13] D. Gosset and D. Nagaj. Quantum 3-SAT is QMA1-complete. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gopalan:2010:FAA

[Gop10] Parikshit Gopalan. A Fourieranalytic approach to Reed– Muller decoding. In IEEE [IEE10], pages 685–694. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Goyal:2012:PRC

[Goy12]  V. Goyal.  Positive results

for concurrently secure computation in the plain model. In IEEE [IEE12], pages 41– 50. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375280.

IEEE Computer Society order number P????.

Goel:2010:OTS

[GP10] Ashish Goel and Ian Post.

One tree suffices: a simultaneous O(1)-approximation for single-sink buy-at-bulk. In IEEE [IEE10], pages 593–600. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Gupta:2013:FDA

[GP13] M. Gupta and R. Peng. Fully dynamic (1 + e)-approximate matchings. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Goldberg:2011:CHM

[GPS11] Paul W. Goldberg, Christos H. Papadimitriou, and Rahul Savani. The complexity of the homotopy method, equilibrium selection, and Lemke– Howson solutions. In IEEE [IEE11], pages 67–76. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Goldwasser:2012:HCP

[GR12] S. Goldwasser and G. N. Rothblum. How to compute in the presence of leakage. In IEEE [IEE12], pages 31–40. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375279.

IEEE Computer Society order number P????.

Guruswami:2010:CCS

[GS10] Venkatesan Guruswami and Adam Smith. Codes for computationally simple channels: Explicit constructions with optimal rate. In IEEE [IEE10], pages 723–732. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Guruswami:2011:LHH

[GS11] Venkatesan Guruswami and Ali Kemal Sinop. Lasserre hierarchy, higher eigenvalues, and approximation schemes for graph partitioning and quadratic integer programming with PSD objectives. In IEEE [IEE11], pages 482–491. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Guruswami:2012:FSH

[GS12] V. Guruswami and A. K.

Sinop. Faster SDP hierarchy solvers for local rounding algorithms. In IEEE [IEE12], pages 197–206. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/

                                        stamp.jsp?arnumber=6375297.

IEEE Computer Society order number P????.

Gabow:2013:AAB

[GS13a] H. N. Gabow and P. Sankowski. Algebraic algorithms for Bmatching, shortest undirected paths, and F -factors. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gaspers:2013:SBB

[GS13b] S. Gaspers and S. Szeider. Strong backdoors to bounded treewidth SAT. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Gharan:2011:RRA

[GSS11] Shayan Oveis Gharan, Amin Saberi, and Mohit Singh. A randomized rounding approach to the traveling salesman problem. In IEEE [IEE11], pages 550–559. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Gharan:2012:AEP

[GT12a] S. O. Gharan and L. Trevisan. Approximating the expansion profile and almost optimal local graph clustering. In IEEE [IEE12], pages 187–196. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375296.

IEEE Computer Society order number P????.

Goel:2012:MOE

[GT12b] G. Goel and P. Tripathi. Matching with our eyes closed. In IEEE [IEE12], pages 718– 727. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375351.

IEEE Computer Society order number P????.

Grandoni:2012:IDS

[GW12] F. Grandoni and V. V. Williams.

Improved distance sensitivity oracles via fast singlesource replacement paths. In IEEE [IEE12], pages 748– 757. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375354.

IEEE Computer Society order number P????.

Guruswami:2013:PCS

[GX13] V. Guruswami and P. Xia. Polar codes: Speed of polarization and polynomial gap to capacity. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Hertli:2011:SFS

[Her11] Timon Hertli. 3-SAT faster and simpler — unique-SAT bounds for PPSZ hold in general. In IEEE [IEE11], pages 277–284. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Huang:2012:EMS

[HK12] Zhiyi Huang and S. Kannan. The exponential mechanism for social welfare: Private, truthful, and nearly optimal. In IEEE [IEE12], pages 140–149. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375291.

IEEE Computer Society order number P????.

Huang:2011:AGS

[HKK11] Zhiyi Huang, Sampath Kannan, and Sanjeev Khanna. Algorithms for the generalized sorting problem. In IEEE [IEE11], pages 738–747. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Henzinger:2013:DAA

[HKN13] M. Henzinger, S. Krinninger, and D. Nanongkai. Dynamic approximate all-pairs shortest paths: Breaking the O(mn) barrier and derandomization. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Hatami:2013:EDT

[HL13] H. Hatami and S. Lovett. Estimating the distance from testable affine-invariant properties. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Hajiaghayi:2013:ONW

[HLP13] M. T. Hajiaghayi, V. Liaghat, and D. Panigrahi. Online nodeweighted Steiner forest and extensions via disk paintings. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Harrow:2010:ETP

[HM10] Aram W. Harrow and Ashley Montanaro. An efficient test for product states with applications to quantum Merlin– Arthur games. In IEEE [IEE10], pages 633–642. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Haitner:2011:CFC

[HO11] Iftach Haitner and Eran Omri. Coin flipping with constant bias implies one-way functions. In IEEE [IEE11], pages 110–119. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Har-Peled:2012:RHR

[HPK12] S. Har-Peled and N. Kumar.

Down the rabbit hole: Robust proximity search and density estimation in sublinear space. In IEEE [IEE12], pages 430–439. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375321.

IEEE Computer Society order number P????.

Har-Peled:2013:AMD

[HPK13] S. Har-Peled and N. Kumar. Approximating minimization diagrams and generalized proximity search. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Hardt:2010:MWM

[HR10] Moritz Hardt and Guy N.

Rothblum. A multiplicative weights mechanism for privacypreserving data analysis. In IEEE [IEE10], pages 61–70. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Holenstein:2012:CPG

[HS12] T. Holenstein and M. Sinha.

Constructing a pseudorandom generator requires an almost linear number of calls. In IEEE [IEE12], pages 698– 707. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375349.

IEEE Computer Society order number P????.

Harris:2013:MTF

[HS13] D. G. Harris and A. Srinivasan. The Moser–Tardos framework with partial resampling. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Haeupler:2010:NCA

[HSS10] Bernhard Haeupler, Barna Saha, and Aravind Srinivasan. New constructive aspects of the Lovász local lemma. In IEEE [IEE10], pages 397–406. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Haramaty:2011:OTM

[HSS11] Elad Haramaty, Amir Shpilka, and Madhu Sudan. Optimal testing of multivariate polynomials over small prime fields. In IEEE [IEE11], pages 629–637. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

IEEE:2010:PIA

[IEE10] IEEE, editor. Proceedings of the 2010 IEEE 51st Annual Symposium on Foundations of Computer Science: 23–26 October 2010, Las Vegas, Nevada, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

IEEE:2011:PIA

[IEE11] IEEE, editor. Proceedings of the 2011 IEEE 52nd Annual Symposium on Foundations of Computer Science: 22–25 October 2011, Palm Springs, California, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=

                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

IEEE:2012:PIA

[IEE12] IEEE, editor. Proceedings of the 2012 IEEE 53rd Annual Symposium on Foundations of Computer Science: 20–23 October 2012, Hyatt Regency, New Brunswick, New Jersey, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS12/;   http://theory.
                                        stanford.edu/~tim/focs12/.

IEEE Computer Society order number P????.

IEEE:2013:PIA

[IEE13]  IEEE, editor.  Proceedings

of the 2012 IEEE 54th Annual Symposium on Foundations of Computer Science: 26–29 October 2013, Berkeley, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Isaksson:2010:GMQ

[IKM10] Marcus Isaksson, Guy Kindler, and Elchanan Mossel. The geometry of manipulation: a quantitative proof of the Gibbard– Satterthwaite theorem. In IEEE [IEE10], pages 319–328. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Kawarabayashi:2010:STM

[iKR10] Ken ichi Kawarabayashi and Bruce Reed. A separator theorem in minor-closed classes. In IEEE [IEE10], pages 153–162. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Kawarabayashi:2011:GMA

[iKRW11] Ken ichi Kawarabayashi, Bruce Reed, and Paul Wollan. The graph minor algorithm with parity conditions. In IEEE [IEE11], pages 27–36. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kawarabayashi:2011:MKW

[iKT11] Ken ichi Kawarabayashi and Mikkel Thorup. The minimum k-way cut of bounded size is fixed-parameter tractable. In IEEE [IEE11], pages 160–169. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Impagliazzo:2012:PS

[IMZ12] R. Impagliazzo, R. Meka, and D. Zuckerman. Pseudorandomness from shrinkage. In IEEE [IEE12], pages 111– 119. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375288.

IEEE Computer Society order number P????.

Impagliazzo:2013:SAS

[IPS13] R. Impagliazzo, R. Paturi, and S. Schneider. A satisfiability algorithm for sparse depth two threshold circuits. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Indyk:2011:PAS

[IPW11] Piotr Indyk, Eric Price, and David P. Woodruff. On the power of adaptivity in sparse recovery. In IEEE [IEE11], pages 285–294. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Ito:2012:MPI

[IV12]  T. Ito and T. Vidick.  A

multi-prover interactive proof for NEXP sound against entangled provers. In IEEE [IEE12], pages 243–252. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375302.

IEEE Computer Society order number P????.

Izumi:2012:NDC

[IW12] T. Izumi and T. Wadayama. A new direction for counting perfect matchings. In IEEE [IEE12], pages 591–598. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375338.

IEEE Computer Society order number P????.

Jain:2012:DPT

[JPY12] R. Jain, A. Pereszlenyi, and Penghui Yao. A direct product theorem for the twoparty bounded-round publiccoin communication complexity. In IEEE [IEE12], pages 167–176. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375294.

IEEE Computer Society order number P????.

Jha:2011:TRL

[JR11] Madhav Jha and Sofya Raskhodnikova. Testing and reconstruction of Lipschitz functions with applications to data privacy. In IEEE [IEE11], pages 433–442. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Jain:2011:PAA

[JY11] Rahul Jain and Penghui Yao. A parallel approximation algorithm for positive semidefinite programming. In IEEE [IEE11], pages 463–471. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kanade:2011:ER

[Kan11a] Varun Kanade. Evolution with recombination. In IEEE [IEE11], pages 837–846. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kane:2011:SPP

[Kan11b] Daniel M. Kane. A small PRG for polynomial threshold functions of Gaussians. In IEEE [IEE11], pages 257–266. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kane:2012:STP

[Kan12]  D. M. Kane.  A structure

theorem for poorly anticoncentrated Gaussian chaoses and applications to the study of polynomial threshold functions. In IEEE [IEE12], pages 91– 100. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375286.

IEEE Computer Society order number P????.

Kumar:2010:CSN

[KK10] Amit Kumar and Ravindran Kannan. Clustering with spectral norm and the k-means algorithm. In IEEE [IEE10], pages 299–308. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Kawarabayashi:2013:ANM

[KK13] K.-I. Kawarabayashi and Y. Kobayashi. All-or-nothing multicommodity flow problem with bounded fractionality in planar graphs. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Kaufman:2011:NEW

[KL11] Tali Kaufman and Shachar Lovett. New extension of the Weil bound for character sums with applications to coding. In IEEE [IEE11], pages 788–796. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kerenidis:2012:LBI

[KLL^{+}12] I. Kerenidis, S. Laplante, V. Lerays, J. Roland, and D. Xiao. Lower bounds on information complexity via zerocommunication protocols and applications. In IEEE [IEE12], pages 500–509. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375328.

IEEE Computer Society order number P????.

Kalai:2012:FRS

[KLR12] Y. T. Kalai, A. Lewko, and A. Rao. Formulas resilient to short-circuit errors. In IEEE [IEE12], pages 490– 499. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375327.

IEEE Computer Society order number P????.

Kolla:2011:HPU

[KMM11] Alexandra Kolla, Konstantin Makarychev, and Yury Makarychev. How to play unique games against a semi-random adversary: Study of semi-random models of unique games. In IEEE [IEE11], pages 443–452. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Koutis:2010:AOS

[KMP10] Ioannis Koutis, Gary L. Miller, and Richard Peng. Approaching optimality for solving SDD linear systems. In IEEE [IEE10], pages 235–244. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Koutis:2011:NML

[KMP11] Ioannis Koutis, Gary L. Miller, and Richard Peng. A nearlym log n time solver for SDD linear systems. In IEEE [IEE11], pages 590–598. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kopelowitz:2012:LIG

[Kop12] T. Kopelowitz. On-line indexing for general alphabets via predecessor queries on subsets of an ordered list. In IEEE [IEE12], pages 283–292. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375306.

IEEE Computer Society order number P????.

Komargodski:2013:IAC

[KRT13] I. Komargodski, R. Raz, and A. Tal. Improved averagecase lower bounds for DeMorgan formula size. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/

FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Khanna:2011:DCC

[KS11a] Sanjeev Khanna and Madhu Sudan. Delays and the capacity of continuous-time channels. In IEEE [IEE11], pages 758–767. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Khot:2011:TPO

[KS11b] Subhash Khot and Muli Safra.

A two prover one round game with strong soundness. In IEEE [IEE11], pages 648–657. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Khot:2012:HFI

[KS12] S. Khot and R. Saket. Hardness of finding independent sets in almost q-colorable graphs. In IEEE [IEE12], pages 380– 389. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375316.

IEEE Computer Society order number P????.

Kempe:2010:FTA

[KSM10] David Kempe, Mahyar Salek, and Cristopher Moore. Frugal and truthful auctions for vertex covers, flows and cuts. In IEEE [IEE10], pages 745–754. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Konemann:2013:LLA

[KSS13] J. Konemann, S. Sadeghian, and L. Sanita. An LMP O(log n)- approximation algorithm for node weighted prize collecting Steiner tree. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Kawarabayashi:2011:MWC

[KT11] K. Kawarabayashi and M. Thorup. The minimum k-way cut of bounded size is fixedparameter tractable. In IEEE [IEE11], pages 160–169. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=

00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Kawarabayashi:2012:CCC

[KT12] K. Kawarabayashi and M. Thorup. Combinatorial coloring of 3-colorable graphs. In IEEE [IEE12], pages 68–75. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375283.

IEEE Computer Society order number P????.

Kratsch:2012:RSI

[KW12] S. Kratsch and M. Wahlstrom.

Representative sets and irrelevant vertices: New tools for kernelization. In IEEE [IEE12], pages 450–459. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375323.

IEEE Computer Society order number P????.

Larsen:2011:RSG

[Lar11] Kasper Green Larsen. On range searching in the group model and combinatorial discrepancy. In IEEE [IEE11], pages 542–549. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Larsen:2012:HCP

[Lar12] K. G. Larsen. Higher cell probe lower bounds for evaluating polynomials. In IEEE [IEE12], pages 293–301. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375307.

IEEE Computer Society order number P????.

Li:2011:MEU

[LD11] Jian Li and Amol Deshpande. Maximizing expected utility for stochastic combinatorial optimization problems. In IEEE [IEE11], pages 797–806. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

LeGall:2012:FAR

[Le 12] F. Le Gall. Faster algorithms for rectangular matrix multiplication. In IEEE [IEE12], pages 514–523. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375330.

IEEE Computer Society order number P????.

Levin:2012:RS

[Lev12] L. A. Levin. Rarity for semimeasures. In IEEE [IEE12], pages 510–513. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375329.

IEEE Computer Society order number P????.

Li:2012:NME

[Li12]  Xin Li.  Non-malleable ex-

tractors, two-source extractors and privacy amplification. In IEEE [IEE12], pages 688–697. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375348.

IEEE Computer Society order number P????.

Li:2013:ECN

[Li13]  Xin Li.  Extractors for a

constant number of independent sources with polylogarithmic min-entropy. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Lovett:2012:CDM

[LM12] S. Lovett and R. Meka. Constructive discrepancy minimization by walking on the edges. In IEEE [IEE12], pages 61– 67. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375282.

IEEE Computer Society order number P????.

Li:2013:IRS

[LMP13] Mu Li, G. L. Miller, and R. Peng. Iterative row sampling. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Lee:2011:QQC

[LMR^{+}11] Troy Lee, Rajat Mittal, Ben W. Reichardt, Robert Spalek, and Mario Szegedy. Quantum query complexity of state conversion. In IEEE [IEE11], pages 344–353. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Lovett:2010:PGC

[LMS10] Shachar Lovett, Partha Mukhopadhyay, and Amir Shpilka. Pseudorandom generators for CC0[p] and the Fourier spectrum of low-degree polynomials over finite fields. In IEEE [IEE10], pages 695–704. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Lacki:2012:SSA

[LNSWN12] J. Lacki, Y. Nussbaum, P. Sankowski, and C. Wulff-Nilsen. Single source–all sinks max flows in planar digraphs. In IEEE [IEE12], pages 599–608. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375339.

IEEE Computer Society order number P????.

Lovett:2010:LBD

[LP10] Shachar Lovett and Ely Porat. A lower bound for dynamic approximate membership data structures. In IEEE [IEE10], pages 797–804. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Lazarus:2012:HTS

[LR12] F. Lazarus and J. Rivaud.

On the homotopy test on surfaces. In IEEE [IEE12], pages 440–449. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375322.

IEEE Computer Society order number P????.

Louis:2013:CAV

[LRV13] A. Louis, P. Raghavendra, and S. Vempala. The complexity of approximating vertex expansion. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Lee:2013:EAC

[LS13] Yin Tat Lee and A. Sidford. Efficient accelerated coordinate descent methods and faster algorithms for solving linear systems. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Leme:2010:PBN

[LT10] Renato Paes Leme and Eva Tardos. Pure and Bayes–Nash price of anarchy for generalized second price auction. In IEEE [IEE10], pages 735–744. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Lewko:2010:IPR

[LW10] Allison Lewko and Brent Waters.

On the insecurity of parallel repetition for leakage resilience. In IEEE [IEE10], pages 521–530. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Madry:2010:FAA

[Mad10] Aleksander Madry. Fast approximation algorithms for cut-based problems in undirected graphs. In IEEE [IEE10], pages 245–254. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Madry:2013:NCP

[Mad13] A. Madry. Navigating central path with electrical flows: From flows to matchings, and back. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Meka:2012:PCS

[Mek12] R. Meka. A PTAS for computing the supremum of Gaussian processes. In IEEE [IEE12], pages 217–222. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375299.

IEEE Computer Society order number P????.

Makarychev:2010:MEO

[MM10] Konstantin Makarychev and Yury Makarychev. Metric extension operators, vertex sparsifiers and Lipschitz extendability. In IEEE [IEE10], pages 255–264. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

McGregor:2010:LTP

[MMP^{+}10] Andrew McGregor, Ilya Mironov, Toniann Pitassi, Omer Reingold, Kunal Talwar, and Salil Vadhan. The limits of twoparty differential privacy. In IEEE [IEE10], pages 81–90. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Mehta:2012:OMS

[MP12] A. Mehta and D. Panigrahi. Online matching with stochastic rewards. In IEEE [IEE12], pages 728–737. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375352.

IEEE Computer Society order number P????.

Maji:2010:CCC

[MPS10] Hemanta K. Maji, Manoj Prabhakaran, and Amit Sahai. On the computational complexity of coin flipping. In IEEE [IEE10], pages 613–622. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Moemke:2011:AGT

[MS11a] Tobias Moemke and Ola Svensson. Approximating graphic TSP by matchings. In IEEE [IEE11], pages 560–569. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Momke:2011:AGT

[MS11b] T. Momke and O. Svensson. Approximating graphic TSP by matchings. In IEEE [IEE11], pages 560–569. ISBN 0-7695-4571-8, 1-4577-1843-X.

ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Moitra:2013:PTA

[MS13] A. Moitra and M. Saks. A polynomial time algorithm for lossy population recovery. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Marcus:2013:IFB

[MSS13] A. Marcus, D. A. Spielman, and N. Srivastava. Interlacing families I: Bipartite Ramanujan graphs of all degrees. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Mulmuley:2012:GCT

[Mul12] K. D. Mulmuley. Geometric complexity theory v: Equivalence between blackbox derandomization of polynomial identity testing and derandomization of Noether’s Normalization Lemma. In IEEE [IEE12], pages 629–638. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375342.

IEEE Computer Society order number P????.

Moitra:2010:SPL

[MV10] Ankur Moitra and Gregory Valiant. Settling the polynomial learnability of mixtures of Gaussians. In IEEE [IEE10], pages 93–102. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Nelson:2013:OFN

[NN13] J. Nelson and H. L. Nguyen.

OSNAP: Faster numerical linear algebra algorithms via sparser subspace embeddings. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Newman:2012:BTP

[NNN12] A. Newman, O. Neiman, and A. Nikolov. Beck’s Three Permutations Conjecture: A counterexample and some consequences. In IEEE [IEE12], pages 253–262. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375303.

IEEE Computer Society order number P????.

Naor:2011:ONW

[NPS11] Joseph (Seffi) Naor, Debmalya Panigrahi, and Mohit Singh. Online node-weighted Steiner tree and related problems. In IEEE [IEE11], pages 210–219. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Potechin:2010:BMS

[Pot10] Aaron Potechin. Bounds on monotone switching networks for directed connectivity. In IEEE [IEE10], pages 553–562. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Patrascu:2010:DOB

[PR10] Mihai Patrascu and Liam Roditty. Distance oracles beyond the Thorup–Zwick bound. In IEEE [IEE10], pages 815–823. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.

org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Patrascu:2012:NID

[PRT12] M. Patrascu, L. Roditty, and M. Thorup. A new infinity of distance oracles for sparse graphs. In IEEE [IEE12], pages 738–747. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375353.

IEEE Computer Society order number P????.

Pruhs:2011:GCA

[Pru11] Kirk Pruhs. Green computing algorithmics. In IEEE [IEE11], pages 3–4. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Poloczek:2012:RGA

[PS12] M. Poloczek and M. Szegedy.

Randomized greedy algorithms for the maximum matching problem with new analysis. In IEEE [IEE12], pages 708– 717. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375350.

IEEE Computer Society order number P????.

Peres:2010:APS

[PSSZ10] Yuval Peres, Dmitry Sotnikov, Benny Sudakov, and Uri Zwick. All-pairs shortest paths in O(n^{2}) time with high probability. In IEEE [IEE10], pages 663–672. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Pagh:2013:HAS

[PSW13] R. Pagh, G. Segev, and U. Wieder. How to approximate a set without knowing its size in advance. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Panigrahy:2010:LBN

[PTW10] Rina Panigrahy, Kunal Talwar, and Udi Wieder. Lower bounds on near neighbor search via metric expansion. In IEEE [IEE10], pages 805–814. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Price:2011:ASR

[PW11a] E. Price and D. P. Woodruff. (1+ ϵ)-approximate sparse recovery. In IEEE [IEE11], pages 295–304. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Price:2011:EAS

[PW11b] Eric Price and David P.

Woodruff. (1 + eps)-approximate sparse recovery. In IEEE [IEE11], pages 295–304. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Rubin:2010:IBG

[RKS10] Natan Rubin, Haim Kaplan, and Micha Sharir. Improved bounds for geometric permutations. In IEEE [IEE10], pages 355–364. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Rossman:2010:MCC

[Ros10a]  B. Rossman.  The mono-

tone complexity of k-clique on random graphs. In IEEE [IEE10], pages 193–201. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Rossman:2010:MCK

[Ros10b] Benjamin Rossman. The monotone complexity of k-clique on random graphs. In IEEE [IEE10], pages 193–201. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Rothvoss:2013:APW

[Rot13] T. Rothvoss. Approximating bin packing within O(log OPT ∗ OPT) bins. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Restrepo:2011:IMC

[RST^{+}11] Ricardo Restrepo, Jinwoo Shin, Prasad Tetali, Eric Vigoda, and Linji Yang. Improved mixing condition on the grid for counting and sampling independent sets. In IEEE [IEE11], pages 140–149. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Ramshaw:2012:WSA

[RT12] L. Ramshaw and R. E. Tarjan. A weight-scaling algorithm for min-cost imperfect matchings in bipartite graphs. In IEEE [IEE12], pages 581– 590. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375337.

IEEE Computer Society order number P????.

Rubin:2013:KDT

[Rub13] N. Rubin. On kinetic Delaunay triangulations: A near quadratic bound for unit speed motions. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Roditty:2011:MWC

[RW11] Liam Roditty and Virginia Vassilevska Williams. Minimum weight cycles and triangles: Equivalences and algorithms. In IEEE [IEE11], pages 180–189. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Santhanam:2010:FPN

[San10]  Rahul Santhanam.  Fight-

ing perebor: New and improved algorithms for formula and QBF satisfiability. In IEEE [IEE10], pages 183–192. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Seguin-Charbonneau:2011:MED

[SCS11] Loic Seguin-Charbonneau and F. Bruce Shepherd. Maximum edge-disjoint paths in planar graphs with congestion 2. In IEEE [IEE11], pages 200–209. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Shaltiel:2011:DAS

[Sha11]  Ronen Shaltiel.  Dispersers

for affine sources with subpolynomial entropy. In IEEE [IEE11], pages 247–256. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Sherman:2013:NMF

[She13] J. Sherman. Nearly maximum flows in nearly linear time. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Sidiropoulos:2010:OSP

[Sid10] Anastasios Sidiropoulos. Optimal stochastic planarization. In IEEE [IEE10], pages 163–170. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Sidiropoulos:2013:NPC

[Sid13] A. Sidiropoulos. Non-positive curvature and the planar embedding conjecture. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.

                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Singer:2010:BFM

[Sin10]  Yaron Singer.  Budget fea-

sible mechanisms. In IEEE [IEE10], pages 765–774. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Sly:2010:CTU

[Sly10] Allan Sly. Computational transition at the uniqueness threshold. In IEEE [IEE10], pages 287–296. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Sohler:2012:AOC

[Soh12] C. Sohler. Almost optimal canonical property testers for satisfiability. In IEEE [IEE12], pages 541–550. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375333.

IEEE Computer Society order number P????.

Saks:2010:ELI

[SS10a] Michael Saks and C. Seshadhri.

Estimating the longest increasing sequence in polylogarithmic time. In IEEE [IEE10], pages 458–467. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Saxena:2010:SGC

[SS10b] Nitin Saxena and C. Seshadhri.

From Sylvester–Gallai configurations to rank bounds: Improved black-box identity test for depth-3 circuits. In IEEE [IEE10], pages 21–29. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Sauerwald:2012:TBR

[SS12a] T. Sauerwald and He Sun. Tight bounds for randomized load balancing on arbitrary network topologies. In IEEE [IEE12], pages 341–350. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375312.

IEEE Computer Society order number P????.

Sly:2012:CHC

[SS12b] A. Sly and N. Sun. The computational hardness of counting in two-spin models on dregular graphs. In IEEE [IEE12], pages 361–369. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375314.

IEEE Computer Society order number P????.

Seeman:2013:ASS

[SS13] L. Seeman and Y. Singer. Adaptive seeding in social networks. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Shah:2011:MAU

[SST11] Devavrat Shah, Jinwoo Shin, and Prasad Tetali. Medium access using queues. In IEEE [IEE11], pages 698–707. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Sinclair:2013:SMA

[SSY13] A. Sinclair, P. Srivastava, and Yitong Yin. Spatial mixing and approximation algorithms for graphs with bounded connective constant. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Saglam:2013:CCS

[ST13] M. Saglam and G. Tardos. On the communication complexity of sparse set disjointness and exists-equal problems. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Thorup:2013:STF

[Tho13]  M. Thorup.  Simple tabu-

lation, fast expanders, double tabulation, and high independence. In IEEE [IEE13], pages 90–99. ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Tsur:2010:TPS

[TR10] Gilad Tsur and Dana Ron. Testing properties of sparse images. In IEEE [IEE10], pages 468–477. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Tulsiani:2011:QGL

[TW11] Madhur Tulsiani and Julia Wolf. Quadratic Goldreich– Levin theorems. In IEEE [IEE11], pages 619–628. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2011&volume=
                                        00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Tsang:2013:FSS

[TWXZ13] Hing Yin Tsang, Chung Hoi Wong, Ning Xie, and Shengyu Zhang. Fourier sparsity, spectral norm, and the log-rank conjecture. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Thapper:2012:PLP

[TZ12] J. Thapper and S. Zivny. The power of linear programming for valued CSPs. In IEEE [IEE12], pages 669–678. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375346.

IEEE Computer Society order number P????.

Vaikuntanathan:2011:CBN

[Vai11] Vinod Vaikuntanathan. Computing blindfolded: New developments in fully homomorphic encryption. In IEEE [IEE11], pages 5–16. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Valiant:2012:FCS

[Val12] G. Valiant. Finding correlations in subquadratic time, with applications to learning parities and juntas. In IEEE [IEE12], pages 11–20. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375277.

IEEE Computer Society order number P????.

Vegh:2012:CGF

[Veg12] L. A. Vegh. Concave generalized flows with applications to market equilibria. In IEEE [IEE12], pages 150–159. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375292.

IEEE Computer Society order number P????.

Vempala:2010:CRS

[Vem10a] Santosh S. Vempala. Corrigendum: A Random Sampling Algorithm for Learning an Intersection of Halfspaces. In IEEE [IEE10], page 123. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Vempala:2010:LCC

[Vem10b] Santosh S. Vempala. Learning convex concepts from Gaussian distributions with PCA. In IEEE [IEE10], pages 124–130. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Viderman:2013:SLI

[Vid13a] M. Viderman. Strong LTCs with inverse poly-log rate and constant soundness. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Vidick:2013:TPE

[Vid13b] T. Vidick. Three-player entangled XOR games are NPhard to approximate. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

http://dimacs.rutgers.edu/
FOCS13/;   http://theory.
stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Viola:2010:CD

[Vio10] Emanuele Viola. The complexity of distributions. In IEEE [IEE10], pages 202–211. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Viola:2011:ECS

[Vio11a] Emanuele Viola. Extractors for circuit sources. In IEEE [IEE11], pages 220–229. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Viola:2011:RBD

[Vio11b]  Emanuele Viola.  Random-

ness buys depth for approximate counting. In IEEE [IEE11], pages 230–239. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Vishnoi:2012:PAT

[Vis12] N. K. Vishnoi. A permanent approach to the Traveling Salesman Problem. In IEEE [IEE12], pages 76–80. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375284.

IEEE Computer Society order number P????.

Valiant:2011:PLE

[VV11] Gregory Valiant and Paul Valiant. The power of linear estimators. In IEEE [IEE11], pages 403–412. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Wee:2010:BBR

[Wee10] Hoeteck Wee. Black-box, roundefficient secure computation via non-malleability amplification. In IEEE [IEE10], pages 531–540. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Wimmer:2010:ALU

[Wim10] Karl Wimmer. Agnostically learning under permutation invariant distributions. In IEEE [IEE10], pages 113–122. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2010&volume=
00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Wulff-Nilsen:2011:STM

[WN11] Christian Wulff-Nilsen. Separator theorems for minor-free and shallow minor-free graphs with applications. In IEEE [IEE11], pages 37–46. ISBN 0-7695-4571-8, 1-4577-1843-X. ISSN 0272-5428. LCCN QA76 .S95 2011. URL http:/

/opac.ieeecomputersociety.
org/opac?year=2011&volume=
00&catalog=4571&acronym=focs.

IEEE Computer Society order number P4571.

Williams:2010:SEB

[WW10] Virginia Vassilevska Williams and Ryan Williams. Subcubic equivalences between path, matrix and triangle problems. In IEEE [IEE10], pages 645–654. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Weimann:2010:RPF

[WY10] Oren Weimann and Raphael Yuster. Replacement paths via fast matrix multiplication. In IEEE [IEE10], pages 655–662. ISBN 0-7695-4244-1, 1-4244-8525-8. ISSN 0272-5428. LCCN QA76 .S95 2010. URL http:/

                                        /opac.ieeecomputersociety.
                                        org/opac?year=2010&volume=
                                        00&catalog=4244&acronym=focs.

IEEE Computer Society order number P4244.

Wigderson:2012:PRP

[WY12] A. Wigderson and A. Yehudayoff. Population recovery and partial identification. In IEEE [IEE12], pages 390– 399. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6375317.

IEEE Computer Society order number P????.

Wilson:2013:FBS

[WZ13] D. B. Wilson and U. Zwick. A forward-backward single-source shortest paths algorithm. In IEEE [IEE13], page ?? ISBN 0-7695-5135-1. ISSN 0272-5428. LCCN QA76 .S95 2012. URL

                                        http://dimacs.rutgers.edu/
                                        FOCS13/;   http://theory.
                                        stanford.edu/~tim/focs13/.

IEEE Computer Society order number P????.

Zhandry:2012:HCQ

[Zha12] M. Zhandry. How to construct quantum random functions. In IEEE [IEE12], pages 679–687. ISBN 1-4673-4383-8. ISSN 0272-5428. LCCN QA76 .S95 2012. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6375347.

IEEE Computer Society order number P????.